### v1

# A central limit theorem for the stochastic cable equation

Abstract
In this paper, we consider one-dimensional nonlinear stochastic cable equations driven by a multiplicative space-time white noise. Using the Malliavin–Stein’s method, we prove the Gaussian fluctuation for the spatial average of the solution under the total variation distance with mild conditions. We also establish functional central limit theorem.
Keywords: Stochastic cable equation, Central limit theorem, Malliain calculus, Stein’s method 0002020 Mathematics Subject Classification: Primary 60H15; Secondary 60F05, 60H07.
1 Introduction and main results
We consider the nonlinear stochastic cable equation
$\frac{\partial u}{\partial t}=\frac{\beta}{2}\frac{\partial^{2}u}{\partial x^{2}}-\alpha u+\sigma(u)\dot{W}$ (1)
on $[0,T]\times[0,L]$ for some constants $\alpha\in\mathbb{R}$ and $\beta>0$, where $T>0$ is fixed, $\dot{W}$ is a space-time white noise on $[0,T]\times[0,L]$, with initial condition $u_{0}(x)=1$, Neumann, Dirichlet, or periodic boundary conditions. We assume that the coefficient $\sigma\in C^{1}(\mathbb{R})$ satisfies global Lipschitz and linear growth conditions.
Following [W86], this equation has a unique mild solution, which is adapted to the filtration generated by $W$, such that $E[u(t,x)^{2}]<\infty$ and it satisfies the evolution equation
$u(t,x)=\int_{0}^{L}u_{0}(y)G_{t}(x,y)\ \mathrm{d}y+\int_{0}^{t}\int_{0}^{L}G_{t-s}(x,y)\sigma(u(s,y))\ W(\mathrm{d}s,\mathrm{d}y)$ (2)
where in the right hand side the stochastic integral is in the sense of Itô–Walsh, and $G$ is the Green’s function for the cable equation (5).
We assume the following technical condition, which is required for the proof of the central limit theorem.
Assumption 1.
There exists some positive valued measurable function $f_{\sigma}$ such that for any $T>0$, $f_{\sigma}\in L^{1}([0,T])$ and for all $t\in[0,T]$,
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[\sigma(u(t,x))^{2}]\ \mathrm{d}x=f_{\sigma}(t).$
In this paper, we are interested in the asymptotic behavior as $L$ tends to infinity of the quantity
$F_{L}(t):=\frac{1}{L}\int_{0}^{L}\left\{u(t,x)-E[u(t,x)]\right\}\ \mathrm{d}x,$ (3)
where $L\geq 1$, $u(t,x)$ is the solution to (1).
Recall that the total variation distance between two random variables $X$ and $Y$ is defined as
$d_{\text{TV}}(X,Y)=\sup_{B\in\mathcal{B}(\mathbb{R})}\left|P(X\in B)-P(Y\in B)\right|$
where $\mathcal{B}(\mathbb{R})$ denotes the Borel $\sigma$-algebra with respect to the Euclidean topology on $\mathbb{R}$. We also write $d_{\text{TV}}(F,\mathcal{N}(0,1))$ for the total variation distance between the law of $F$ and the standard normal law.
We are now ready to state the first main result of this paper.
Theorem 1.
Suppose that $\operatorname{Var}(F_{L}(t))>0$. Then, for every $t>0$ there exists a real number $c=c(t)>0$ such that for all $L\geq 1$,
$d_{\mathrm{TV}}\left(\frac{F_{L}(t)}{\sqrt{\operatorname{Var}(F_{L}(t))}},\mathcal{N}(0,1)\right)\leq\frac{c}{\sqrt{L}}.$ (4)
We also prove the following functional central limit theorem.
Theorem 2.
Fix $T>0$. Then, as $L\to\infty$,
$\left(\sqrt{L}F_{L}(t)\right)_{t\in[0,T]}\to\left(\int_{0}^{t}e^{-\alpha(t-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}\right)_{t\in[0,T]}$
where $f_{\sigma}(t)$ is the limit in Assumption 1 and $W=\{W_{s}\}_{s\in[0,T]}$ denotes a standard one-dimensional Brownian motion, and the convergence is in law on the space of continuous functions $C([0,T])$.
This kind of asymptotic behavior of spatial averages for stochastic partial differential equations has been studied extensively in recent years. This line of research was initiated by Huang, Nualart, and Viitasaari in [HNV20], where they established quantitative central limit theorem and its functional version for one-dimensional nonlinear stochastic heat equations driven by space-time white noise using Malliavin–Stein’s method (cf. [NP12]). After this result, similar central limit theorems for stochastic partial differential equations have been established under various settings. On the other hand, many of these studies deal with SPDEs whose solutions exhibit spatial stationarity. In our case, as well as in [P22], we consider SPDEs on the interval $[0,L]$, and therefore the solutions are not stationary in space, which requires technical conditions such as Assumption 1 for the proof of the central limit theorem. Moreover, the stochastic cable equation studied in this paper can be seen as a generalization of the stochastic heat equation considered in [P22]. For these reasons, our study can be regarded as a generalization of [P22].
Assumption 1 is verified in Section 6 for the case $\sigma(u)=\sigma_{1}u+\sigma_{0}$, where $\sigma_{1}$ and $\sigma_{0}$ are constants. The argument relies on the Wiener chaos decomposition of the solution to the stochastic cable equation (1). The same method was used in [P22] to establish Assumption 1 for the stochastic heat equation with $\sigma(u)=u$. In [HNV20] and related works, since the solution to the SPDE under consideration is spatially stationary, the limit in Assumption 1 reduces to $E[\sigma(u(t,0))^{2}]$, and thus technical conditions such as those in Assumption 1 are not required.
The nonlinear stochastic cable equation emerges within the framework of mathematical models of neurons, where it governs the propagation of electrical signals along the cylindrical structure of neurons. Neurons, the fundamental units of the nervous system, operate through an intricate interplay of chemical, biological, and electrical processes. In a standard mathematical idealization, neurons are represented as long, thin cylinders, analogous to electrical cables, enabling the application of cable theory. For such a cylindrical structure spanning from $x=0$ to $x=L$, we denote the electrical potential at position $x$ and time $t$ by $u(t,x)$, focusing on variations along the spatial coordinate $x$. This potential $u(t,x)$ is governed by a system of non-linear PDE’s, called the Hodgkin-Huxley equations, but in certain ranges of values of $u$, they are well-approximated by the cable equation:
$\frac{\partial u}{\partial t}=\frac{\beta}{2}\frac{\partial^{2}u}{\partial x^{2}}-\alpha u,$ (5)
where $\beta/2>0$ represents the rate of diffusion within the neuron and $\alpha\in\mathbb{R}$ is the rate at which ions leak across the membrance [KX93]. The surface of the neuron is covered with synapses, through which it receives impulses of current. If the current arriving at $(t,x)$ is denoted by $F(t,x)$, the system satisfies the inhomogeneous PDE:
$\frac{\partial u}{\partial t}=\frac{\beta}{2}\frac{\partial^{2}u}{\partial x^{2}}-\alpha u+F.$
Even when the system is at rest, occasional random impulses may occur, implying that $F$ generally contains a stochastic component. In particular, Walsh studied the case where $F$ is compound Poisson process or space-time white noise [W81]. In [W86], the nonlinear stochastic cable equation was studied for the case $F=\sigma(u)\dot{W}$.
2 Preliminaries
We write $\|Z\|_{k}$ instead of $E[|Z|^{k}]^{1/k}$, for every $Z\in L^{k}(\Omega)$.
2.1 Clark–Ocone formula
Let $\mathcal{H}=L^{2}([0,T]\times\mathbb{R})$. The Gaussian family $\{X(h)\}_{h\in\mathcal{H}}$ formed by the stochastic integrals
$X(h)=\int_{[0,T]\times\mathbb{R}}h(s,x)\ W(\mathrm{d}s,\mathrm{d}x)$
defines an isonormal Gaussian process on the Hilbert space $\mathcal{H}$. In this framework we can develop the Malliavin calculus [N06]. We denote by $D$ the Malliavin derivative operator. Let $\{\mathcal{F}_{s}\}_{s\in[0,T]}$ denote the filtration generated by the space-time white noise $\dot{W}$. We recall the following Clark–Ocone formula:
$F=E[F]+\int_{[0,T]\times\mathbb{R}}E[D_{s,z}F\mid\mathcal{F}_{s}]\ W(\mathrm{d}s,\mathrm{d}z)\mathrm{~{}a.s.},$
valid for every $\mathcal{F}_{T}$measurable random variable $F$ in the Sobolev space $\mathbb{D}^{1,2}$ in the sense of Malliavin derivative. By using Jensen’s inequality for conditional expectations, the above Clark–Ocone formula yields the following Poincaré-type inequality:
$\left|\operatorname{Cov}[F,G]\right|\leq\int_{0}^{T}\int_{\mathbb{R}}\|D_{s,z}F\|_{2}\|D_{s,z}G\|_{2}\ \mathrm{d}z\mathrm{d}s$ (6)
for all $\mathcal{F}_{T}$measurable random variables $F,G\in\mathbb{D}^{1,2}$.
2.2 The Malliavin–Stein’s method
Let us now introduce some preliminary results of Malliavin–Stein’s method [NP12]. Stein’s method is a probabilistic technique to allow one to get some bounds for the distance between two probability measures. Combining Stein’s method and the Malliavin calculus, we can obtain useful estimate to prove quantitative central limit theorems.
Proposition 2.1.
Let $F=\delta(v)$ for some $v\in\operatorname{Dom}(\delta)$ which is the $L^{2}(\Omega)$-domain of the adjoint of the Malliavin derivative operator. Suppose that $F\in\mathbb{D}^{1,2}$ and $E[F^{2}]=1$. Then we have
$d_{\mathrm{TV}}(F,\mathcal{N}(0,1))\leq 2\sqrt{\operatorname{Var}(\langle DF,v\rangle_{\mathcal{H}})}.$
In order to establish the functional CLT in Theorem 2, we also need the following multivariate counterpart of Proposition 2.1.
Proposition 2.2.
Fix $m\geq 2$. Let $F=(F_{1},\dots,F_{m})$ be a random vector such that for every $i=1,\dots,m$, $F_{i}=\delta(v_{i})$ for some $v_{i}\in\operatorname{Dom}(\delta)$ and $F_{i}\in\mathbb{D}^{1,2}$. Let $Z$ be an $m$-dimensional centered Gaussian random vector with covariance matrix $(C_{i,j})_{1\leq i,j\leq m}$. Then, for every $h\in C^{2}(\mathbb{R}^{m})$ that has bounded second partial derivatives,
$\left|E[h(F)]-E[h(Z)]\right|\leq\frac{1}{2}\|h^{\prime\prime}\|_{\infty}\sqrt{\sum_{i,j=1}^{m}E\left[\left|C_{i,j}-\langle DF_{i},v_{j}\rangle_{\mathcal{H}}\right|^{2}\right]},$
where
$\|h^{\prime\prime}\|_{\infty}:=\max_{1\leq i,j\leq m}\sup_{x\in\mathbb{R}^{m}}\left|\frac{\partial^{2}h(x)}{\partial x_{i}\partial x_{j}}\right|.$
2.3 Moments and Malliavin derivative of $u(t,x)$
Recently, Pu proved that the moments of $u(t,x)$ are uniformly bounded for all $L\geq 1$ in the case where $\alpha=0$ and $\beta=1$ [P22]*Lemma 2.3.. Similarly, a corresponding result has been obtained for the case $\alpha\in\mathbb{R}$ and $\beta>0$.
Lemma 2.1.
Let $u(t,x)$ be the solution to (1). Then for all $k\geq 2$,
$\sup_{L\geq 1}\sup_{(t,x)\in[0,T]\times[0,L]}\|u_{n}(t,x)\|_{k}\leq c_{T,k}<\infty.$ (7)
Pu also proved that the moments of the Malliavin derivative of $u(t,x)$ satisfy a Gaussian-type upper bound uniformly over $L\geq 1$ in the case where $\alpha=0$ and $\beta=1$ [P22]*Lemma 2.4.. This result can similarly be extended to the case $\alpha\in\mathbb{R}$ and $\beta>0$.
Lemma 2.2.
Fix $T>0$. Then, for all $k\geq 2$, there exists $C_{T,k}>0$ such that for all $L\geq 1$, $(t,x)\in[0,T]\times[0,L]$, and for almost every $(s,y)\in(0,t)\times\mathbb{R}$,
$\|D_{s,y}u(t,x)\|_{k}\leq\begin{cases}C_{T,k}1_{[0,L]}(y)p_{\beta(t-s)}(x-y)&\mathrm{(Neumann/Dirichlet~{}case)}\\ C_{T,k}1_{[0,L]}(y)G_{t-s}(x-y)&\mathrm{(periodic~{}case)}\end{cases}$
where $p_{t}(z)$ denotes the Gaussian heat kernel, defined in (25). In particular, we have $u(t,x)\in\bigcap_{k\geq 2}\mathbb{D}^{1,k}$ for all $(t,x)\in[0,T]\times[0,L]$.
3 Asymptotic behavior of the covariance
In this section, we analyze the asymptotic behavior of the covariance of the spatial integral of the solution to (1).
Since (26), by the same arguments as [P22]*Lemma 3.2. and Lemma A.4., we obtain the following supporting lemma.
Lemma 3.1.
Fix $T>0$. Denote for $(t,x)\in(0,T]\times[0,L]$,
$\mathcal{I}_{0}(t,x)=\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y,\text{~{}and}\quad\mathcal{I}_{0}(0,x)=1.$
Then
$\sup_{L\geq 1}\sup_{(t,x)\in[0,T]\times[0,L]}\mathcal{I}_{0}(t,x)\leq e^{|\alpha|T},$
and for all $t>0$,
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}\mathcal{I}_{0}(t,x)\ \mathrm{d}x=e^{-\alpha t},$
and for all $t_{1},t_{2}>0$,
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}\mathcal{I}_{0}(t_{1},x)\mathcal{I}_{0}(t_{2},x)\ \mathrm{d}x=e^{-\alpha(t_{1}+t_{2})}.$
The following result provides the asymptotic behavior of the covariance function of the renormalized sequence of processes $F_{L}(t)$ as $L$ tends to infinity.
Proposition 3.1.
For every $t_{1},t_{2}>0$,
$\lim_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]=\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}f_{\sigma}(s)\ \mathrm{d}s,$
where the function $f_{\sigma}$ is defined in Assumption 1.
Proof.
Using the mild form in (2) and Itô’s isometry, we have
$\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$
$=\frac{1}{L}\int_{[0,L]^{2}}\operatorname{Cov}\left[u(t_{1},x),u(t_{2},y)\right]\ \mathrm{d}x\mathrm{d}y$
$=\frac{1}{L}\int_{[0,L]^{2}}\int_{0}^{t_{1}\wedge t_{2}}\int_{0}^{L}G_{t_{1}-s}(x,z)G_{t_{2}-s}(y,z)E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s\mathrm{d}x\mathrm{d}y$
$=\frac{1}{L}\int_{0}^{t_{1}\wedge t_{2}}\int_{0}^{L}\mathcal{I}_{0}(t_{1}-s,z)\mathcal{I}_{0}(t_{2}-s,z)E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s,$
where $\mathcal{I}_{0}$ is defined in Lemma 3.1. Moreover, we obtain
$\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$
$=\frac{1}{L}\int_{0}^{t_{1}\wedge t_{2}}\int_{0}^{L}\left[\mathcal{I}_{0}(t_{1}-s,z)\mathcal{I}_{0}(t_{2}-s,z)-e^{-\alpha(t_{1}+t_{2}-2s)}\right]E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s$
$\quad+\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}\frac{1}{L}\int_{0}^{L}E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s.$
By (7),
$\left|\frac{1}{L}\int_{0}^{t_{1}\wedge t_{2}}\int_{0}^{L}\left[\mathcal{I}_{0}(t_{1}-s,z)\mathcal{I}_{0}(t_{2}-s,z)-e^{-\alpha(t_{1}+t_{2}-2s)}\right]E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s\right|$
$\leq 2M_{\sigma}^{2}(1+c_{T,2}^{2})\int_{0}^{t_{1}\wedge t_{2}}\left|\frac{1}{L}\int_{0}^{L}\mathcal{I}_{0}(t_{1}-s,z)\mathcal{I}_{0}(t_{2}-s,z)\ \mathrm{d}z-e^{-\alpha(t_{1}+t_{2}-2s)}\right|\ \mathrm{d}s,$
where $M_{\sigma}>0$ is the constant in the linear growth condition for $\sigma$. Hence, applying Lemma 3.1 and dominated convergence theorem, we obtain
$\frac{1}{L}\int_{0}^{t_{1}\wedge t_{2}}\int_{0}^{L}\left[\mathcal{I}_{0}(t_{1}-s,z)\mathcal{I}_{0}(t_{2}-s,z)-e^{-\alpha(t_{1}+t_{2}-2s)}\right]E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s$
$\to 0\mathrm{~{}as~{}}L\to\infty.$
Then, by Assumption 1, we have
$\lim_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$
$=\lim_{L\to\infty}\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}\frac{1}{L}\int_{0}^{L}E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s$
$=\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}f_{\sigma}(s)\ \mathrm{d}s.$
4 Proof of Theorem 1
Using stochastic Fubini’s theorem, we have
$F_{L}(t)=\int_{0}^{t}\int_{0}^{L}v_{L,t}(s,y)\ W(\mathrm{d}s,\mathrm{d}y)=\delta(v_{L,t})\mathrm{~{}a.s.},$ (8)
where
$v_{L,t}(s,y)$ (9)
$:=\frac{1}{L}1_{(0,t)}(s)1_{[0,L]}(y)\sigma(u(s,y))\int_{0}^{L}G_{t-s}(x,y)\ \mathrm{d}x$
$=\frac{1}{L}1_{(0,t)}(s)1_{[0,L]}(y)\sigma(u(s,y))\mathcal{I}_{0}(t-s,y),$
and $\delta$ is the adjoint of the Malliavin derivative operator.
We prepare the following proposition for the proof of our main results.
Proposition 4.1.
For every $T>0$, there exists $A_{T}>0$ such that
$\sup_{t,\tau\in[0,T]}\operatorname{Var}\left(\left\langle DF_{L}(t),v_{L,\tau}\right\rangle_{\mathcal{H}}\right)\leq\frac{A_{T}}{L^{3}}\text{~{}for~{}all~{}}L\geq 1.$
Proof.
From Proposition $1.3.2$ of [N06] and (8), we have
$D_{r,z}F_{L}(t)=1_{(0,t)}(r)v_{L,t}(r,z)+1_{(0,t)}(r)\int_{r}^{t}\int_{0}^{L}D_{r,z}v_{L,t}(s,y)\ W(\mathrm{d}s,\mathrm{d}y).$
Hence, using the stochastic Fubini’s theorem, we obtain
$\left\langle DF_{L}(t),v_{L,\tau}\right\rangle_{\mathcal{H}}$
$=\left\langle v_{L,t},v_{L,\tau}\right\rangle_{\mathcal{H}}+\int_{0}^{\tau}\int_{0}^{L}v_{L,\tau}(r,z)\left(\int_{r}^{t}\int_{0}^{L}D_{r,z}v_{L,t}(s,y)\ W(\mathrm{d}s,\mathrm{d}y)\right)\mathrm{d}z\mathrm{d}r$
$=\left\langle v_{L,t},v_{L,\tau}\right\rangle_{\mathcal{H}}+\int_{0}^{t}\int_{0}^{L}\left(\int_{0}^{\tau\wedge s}\int_{0}^{L}v_{L,\tau}(r,z)D_{r,z}v_{L,t}(s,y)\ \mathrm{d}z\mathrm{d}r\right)W(\mathrm{d}s,\mathrm{d}y).$
Therefore,
$\operatorname{Var}\left(\left\langle DF_{L}(t),v_{L,\tau}\right\rangle_{\mathcal{H}}\right)\leq 2(\Phi_{L,t,\tau}^{(1)}+\Phi_{L,t,\tau}^{(2)}),$
where
$\Phi_{L,t,\tau}^{(1)}=\operatorname{Var}\left(\left\langle v_{L,t},v_{L,\tau}\right\rangle_{\mathcal{H}}\right),$
$\Phi_{L,t,\tau}^{(2)}=\operatorname{Var}\left(\int_{0}^{t}\int_{0}^{L}\left(\int_{0}^{\tau\wedge s}\int_{0}^{L}v_{L,\tau}(r,z)D_{r,z}v_{L,t}(s,y)\ \mathrm{d}z\mathrm{d}r\right)W(\mathrm{d}s,\mathrm{d}y)\right).$
From (9),
$\Phi_{L,t,\tau}^{(1)}$
$=\frac{1}{L^{4}}\operatorname{Var}\left(\int_{0}^{t\wedge\tau}\int_{0}^{L}\sigma(u(s,y))^{2}\mathcal{I}_{0}(t-s,y)\mathcal{I}_{0}(\tau-s,y)\ \mathrm{d}y\mathrm{d}s\right)$
$=\frac{1}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\operatorname{Cov}\left[\sigma(u(s_{1},y_{1}))^{2},\sigma(u(s_{2},y_{2}))^{2}\right]$
×$\quad\times\mathcal{I}_{0}(t-s_{1},y_{1})\mathcal{I}_{0}(\tau-s_{1},y_{1})\mathcal{I}_{0}(t-s_{2},y_{2})\mathcal{I}_{0}(\tau-s_{2},y_{2})\ \mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}.$
By Lemma 3.1, Hölder’s inequality and Minkowski inequality, we have
$\Phi_{L,t,\tau}^{(1)}$
$\leq\frac{4e^{4|\alpha|T}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{\mathbb{R}}\|\sigma(u(s_{1},y_{1}))D_{r,z}\sigma(u(s_{1},y_{1}))\|_{2}$
×$\quad\times\|\sigma(u(s_{2},y_{2}))D_{r,z}\sigma(u(s_{2},y_{2}))\|_{2}\ \mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
$\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{\mathbb{R}}\|D_{r,z}\sigma(u(s_{1},y_{1}))\|_{4}\|D_{r,z}\sigma(u(s_{2},y_{2}))\|_{4}$
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative [N06],
$D_{r,z}\sigma(u(s,y))=\sigma^{\prime}(u(s,y))D_{r,z}u(s,y).$
Then, we obtain
$\Phi_{L,t,\tau}^{(1)}$
$\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{\mathbb{R}}\|D_{r,z}u(s_{1},y_{1})\|_{4}\|D_{r,z}u(s_{2},y_{2})\|_{4}$
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
where $K_{\sigma}$ is Lipschitz constant of $\sigma$. In the case of Neumann/Dirichlet boundary conditions, by Lemma 2.2, we have
$\Phi_{L,t,\tau}^{(1)}$
$\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{\mathbb{R}}p_{\beta(s_{1}-r)}(y_{1}-z)p_{\beta(s_{2}-r)}(y_{2}-z)$
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
$=\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}p_{\beta(s_{1}+s_{2}-2r)}(y_{1}-y_{2})$
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
where the equality holds by the semigroup property of Gaussian heat kernel $p_{t}(z)$. By using the identity
$\int_{\mathbb{R}}p_{\beta(s_{1}+s_{2}-2r)}(y_{1})\mathrm{d}y_{1}=1,$ (10)
we obtain
$\Phi_{L,t,\tau}^{(1)}$
$\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]}\int_{\mathbb{R}}\int_{0}^{s_{1}\wedge s_{2}}p_{\beta(s_{1}+s_{2}-2r)}(y_{1})$
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
$=\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}\int_{[0,t\wedge\tau]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\ \mathrm{d}r\mathrm{d}s_{1}\mathrm{d}s_{2}$
$\leq\frac{4T^{3}e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}.$
In the case of periodic boundary conditions, from Lemma 2.2, we have
$\Phi_{L,t,\tau}^{(1)}$
$\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{0}^{L}G_{s_{1}-r}(y_{1},z)G_{s_{2}-r}(y_{2},r)$
×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
$=\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}\int_{[0,t\wedge\tau]^{2}}\int_{0}^{s_{1}\wedge s_{2}}e^{-\alpha(s_{1}+s_{2}-r)}\mathrm{d}r\mathrm{d}s_{1}\mathrm{d}s_{2}$
$\leq\frac{4T^{3}e^{6|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}},$
where the equality follows from (28).
We proceed to estimate $\Phi_{L,t,\tau}^{(2)}$. By Itô’s isometry, we have
$\Phi_{L,t,\tau}^{(2)}$
$=\int_{0}^{t}\int_{0}^{L}\left\|\int_{0}^{\tau\wedge s}\int_{0}^{L}v_{L,\tau}(r,z)D_{r,z}v_{L,t}(s,y)\right\|_{2}^{2}\ \mathrm{d}y\mathrm{d}s$
$=\frac{1}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\mathcal{I}_{0}(\tau-r_{1},z_{1})\mathcal{I}_{0}(\tau-r_{2},z_{2})\mathcal{I}_{0}^{2}(t-s,y)$
×$\quad\times E[\sigma(u(r_{1},z_{1}))(D_{r_{1},z_{1}}\sigma(u(s,y)))\sigma(u(r_{2},z_{2}))(D_{r_{2},z_{2}}\sigma(u(s,y)))]\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s.$
By Lemma 3.1, Hölder’s inequality and Minkowski inequality,
$\Phi_{L,t,\tau}^{(2)}$
$\leq\frac{e^{4|\alpha|T}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|\sigma(u(r_{1},z_{1}))\|_{4}\|D_{r_{1},z_{1}}\sigma(u(s,y))\|_{4}$
×$\quad\times\|\sigma(u(r_{2},z_{2}))\|_{4}\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|D_{r_{1},z_{1}}\sigma(u(s,y))\|_{4}$
×$\quad\times\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative [N06], we obtain
$\Phi_{L,t,\tau}^{(2)}$
$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|\sigma^{\prime}(u(s,y))D_{r_{1},z_{1}}u(s,y)\|_{4}$
×$\quad\times\|\sigma^{\prime}(u(s,y))D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|D_{r_{1},z_{1}}u(s,y)\|_{4}$
×$\quad\times\|D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
where $K_{\sigma}$ is Lipschitz constant of $\sigma$. In the case of Neumann/Dirichlet boundary conditions, applying Lemma 2.2, we have
$\Phi_{L,t,\tau}^{(2)}$
$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}$
×$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\int_{\mathbb{R}}$
×$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}y\mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
$=\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}$
×$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s,$
where the equality holds by the semigroup property of Gaussian heat kernel $p_{t}(z)$. By using the identity (10), we obtain
$\Phi_{L,t,\tau}^{(2)}$
$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}\int_{0}^{L}\int_{\mathbb{R}}$
×$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
$=\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
$\leq\frac{T^{3}e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}$
In the case of periodic boundary conditions, applying Lemma 2.2 and identity (28), we have
$\Phi_{L,t,\tau}^{(2)}$
$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}$
×$\quad\times G_{s-r_{1}}(y,z_{1})G_{s-r_{2}}(y,z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
$=\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}e^{-\alpha(2s-r_{1}-r_{2})}\ \mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
$\leq\frac{T^{3}e^{6|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}.$
This completes the proof. ∎
We are now ready to prove Theorem 1.
Proof of Theorem 1.
Applying Proposition 4.1 with $t=\tau$, we have
$\sup_{t\in[0,T]}\operatorname{Var}\left(\left\langle DF_{L}(t),v_{L,t}\right\rangle_{\mathcal{H}}\right)\leq\frac{A_{T}}{L^{3}}\text{~{}for~{}all~{}}L\geq 1$
for all $L\geq 1$. Then, by using (8) and Proposition 2.1,
$d_{\mathrm{TV}}\left(\frac{F_{L}(t)}{\sqrt{\operatorname{Var}(F_{L}(t))}},\mathcal{N}(0,1)\right)$ (11)
$\leq 2\sqrt{\operatorname{Var}\left(\left\langle\frac{DF_{L}(t)}{\sqrt{\operatorname{Var}(F_{L}(t))}},\frac{v_{L,t}}{\sqrt{\operatorname{Var}(F_{L}(t))}}\right\rangle_{\mathcal{H}}\right)}$
$\leq\frac{2\sqrt{A_{t}}}{L^{3/2}\operatorname{Var}(F_{L}(t))}$
uniformly for all $t\in[0,T]$ and $L\geq 1$. Proposition 3.1 ensures that
$\operatorname{Var}(F_{L}(t))\sim\frac{1}{L}\int_{0}^{t}e^{-2\alpha(t-s)}f_{\sigma}(s)\ \mathrm{d}s$
as $L\to\infty$. This together with (11) implies (4). ∎
5 Proof of Theorem 2
We prepare the following proposition for the proof of Theorem 2.
Proposition 5.1.
For every $T>0$ and $k\geq 2$, there exists $A_{T,k}>0$ such that for all $t_{1},t_{2}\in[0,T]$,
$\|F_{L}(t_{2})-F_{L}(t_{1})\|_{k}\leq A_{T,k}|t_{2}-t_{1}|^{1/2}L^{-1/2}$ (12)
uniformly for all $L\geq 1$.
Proof.
From (3),
$F_{L}(t)=\frac{1}{L}\int_{0}^{t}\int_{0}^{L}\mathcal{I}_{0}(t-s,y)\sigma(u(s,y))\ W(\mathrm{d}s,\mathrm{d}y).$ (13)
We assume that $t_{1}\leq t_{2}$. Applying Burkholder’s inequality, for all $k\geq 2$,
$\|F_{L}(t_{2})-F_{L}(t_{1})\|_{k}^{2}$
$\leq\frac{2z_{k}^{2}}{L^{2}}\left(\int_{t_{1}}^{t_{2}}\int_{0}^{L}\mathcal{I}_{0}^{2}(t_{2}-s,y)\|\sigma(u(s,y))\|_{k}^{2}\ \mathrm{d}y\mathrm{d}s\right.$
$\quad+\left.\int_{0}^{t_{1}}\int_{0}^{L}(\mathcal{I}_{0}(t_{2}-s,y)-\mathcal{I}_{0}(t_{1}-s,y))^{2}\|\sigma(u(s,y))\|_{k}^{2}\ \mathrm{d}y\mathrm{d}s\right),$
where $z_{k}$ is the constant in Burkholder’s inequality. By using (7), Minkowski’s inequality and the change of variables,
$\|F_{L}(t_{2})-F_{L}(t_{1})\|_{k}^{2}$ (14)
$\leq\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L^{2}}\left(\int_{t_{1}}^{t_{2}}\int_{0}^{L}\mathcal{I}_{0}^{2}(t_{2}-s,y)\ \mathrm{d}y\mathrm{d}s\right.$
$\quad+\left.\int_{0}^{t_{1}}\int_{0}^{L}(\mathcal{I}_{0}(t_{2}-s,y)-\mathcal{I}_{0}(t_{1}-s,y))^{2}\ \mathrm{d}y\mathrm{d}s\right)$
$=\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L^{2}}\left(\int_{t_{1}}^{t_{2}}\int_{0}^{L}\mathcal{I}_{0}^{2}(t_{2}-s,y)\ \mathrm{d}y\mathrm{d}s\right.$
$\quad+\left.\int_{0}^{t_{1}}\int_{0}^{L}(\mathcal{I}_{0}(t_{2}-t_{1}+s,y)-\mathcal{I}_{0}(s,y))^{2}\ \mathrm{d}y\mathrm{d}s\right),$
where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. In the case of Neumann/periodic boundary conditions, using $\mathcal{I}_{0}(t,y)=e^{-\alpha t}$(see Lemma 3.1 and (28) in Lemma A.1),
$\|F_{L}(t_{2})-F_{L}(t_{1})\|_{k}^{2}$
$\leq\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L^{2}}\left(\int_{t_{1}}^{t_{2}}\int_{0}^{L}e^{-2\alpha(t_{2}-s)}\ \mathrm{d}y\mathrm{d}s\right.$
$\quad+\left.\int_{0}^{t_{1}}\int_{0}^{L}(e^{-\alpha(t_{2}-t_{1}+s)}-e^{-\alpha s})^{2}\ \mathrm{d}y\mathrm{d}s\right).$
From the following inequality
$|e^{-\alpha(t_{2}-t_{1})}-1|\leq|\alpha|e^{|\alpha|T}|t_{2}-t_{1}|\quad\mathrm{for}\ t_{1},t_{2}\in[0,T],$
we have
$\|F_{L}(t_{2})-F_{L}(t_{1})\|_{k}^{2}$
$\leq\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L^{2}}\left(\int_{t_{1}}^{t_{2}}\int_{0}^{L}e^{2|\alpha|T}\ \mathrm{d}y\mathrm{d}s\right.$
$\quad+\left.\int_{0}^{t_{1}}\int_{0}^{L}e^{4|\alpha|T}|\alpha|^{2}|t_{2}-t_{1}|^{2}\ \mathrm{d}y\mathrm{d}s\right)$
$\leq\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L}\left(e^{2|\alpha|T}|t_{2}-t_{1}|+Te^{4|\alpha|T}|\alpha|^{2}|t_{2}-t_{1}|^{2}\right)$
$\leq\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L}\left(e^{2|\alpha|T}+T^{2}e^{4|\alpha|T}|\alpha|^{2}\right)|t_{2}-t_{1}|.$
In the case of Dirichlet boundary conditions, using the representation of Dirichlet heat kernel (27),
$\mathcal{I}_{0}(t_{2}-t_{1}+s,y)-\mathcal{I}_{0}(s,y)$
$=\frac{2}{L}\sum_{n=1}^{\infty}\sin\left(\frac{n\pi y}{L}\right)\int_{0}^{L}\sin\left(\frac{n\pi z}{L}\right)\ \mathrm{d}z\left[e^{-\alpha(t_{2}-t_{1}+s)}e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1}+s)}{2L^{2}}}-e^{-\alpha s}e^{-\frac{n^{2}\pi^{2}s}{2L^{2}}}\right]$
$=\frac{2e^{-\alpha s}}{L}\sum_{n=1}^{\infty}\sin\left(\frac{n\pi y}{L}\right)\frac{1-\cos(n\pi)}{n\pi}e^{-\frac{n^{2}\pi^{2}s}{2L^{2}}}\left[e^{-\alpha(t_{2}-t_{1})}e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right].$
Then, applying the $L^{2}([0,L])$-orthogonality of the functions $\{y\mapsto\sin(n\pi y/L)\}_{n=1}^{\infty}$, we obtain
$\int_{0}^{t_{1}}\int_{0}^{L}(\mathcal{I}_{0}(t_{2}-t_{1}+s,y)-\mathcal{I}_{0}(s,y))^{2}\ \mathrm{d}y\mathrm{d}s$
$=4\int_{0}^{t_{1}}e^{-2\alpha s}\sum_{n=1}^{\infty}\int_{0}^{L}\sin^{2}\left(\frac{n\pi y}{L}\right)\ \mathrm{d}y\left(\frac{1-\cos(n\pi)}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\left[e^{-\alpha(t_{2}-t_{1})}e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right]^{2}\ \mathrm{d}s$
$=2L\int_{0}^{t_{1}}e^{-2\alpha s}\sum_{n=1}^{\infty}\left(\frac{1-\cos(n\pi)}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\left[e^{-\alpha(t_{2}-t_{1})}e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right]^{2}\ \mathrm{d}s$
$\leq 8L\int_{0}^{t_{1}}e^{-2\alpha s}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\left[e^{-\alpha(t_{2}-t_{1})}e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right]^{2}\ \mathrm{d}s.$
By using the following inequality
$\left|e^{-\alpha(t_{2}-t_{1})}e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|$
$\leq\left|e^{-\alpha(t_{2}-t_{1})}e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}\right|+\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|$
$=e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}\left|e^{-\alpha(t_{2}-t_{1})}-1\right|+\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|$
$\leq|\alpha|e^{|\alpha|T}|t_{2}-t_{1}|+\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|\quad\mathrm{for}\ t_{1},t_{2}\in[0,T],$
we have
$\int_{0}^{t_{1}}\int_{0}^{L}(\mathcal{I}_{0}(t_{2}-t_{1}+s,y)-\mathcal{I}_{0}(s,y))^{2}\ \mathrm{d}y\mathrm{d}s$
$\leq 16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\int_{0}^{t_{1}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\ \mathrm{d}s\left(|\alpha|^{2}e^{2|\alpha|T}|t_{2}-t_{1}|^{2}+\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}\right)$
$=\frac{16}{6}L|\alpha|^{2}Te^{4|\alpha|T}|t_{2}-t_{1}|^{2}+16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\int_{0}^{t_{1}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\ \mathrm{d}s\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}.$
Using the fact that $1-e^{-x}\leq 1\wedge x$ for all $x\geq 0$,
$16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\int_{0}^{t_{1}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\ \mathrm{d}s\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}$
$=16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\frac{L^{2}}{n^{2}\pi^{2}}\left(1-e^{-\frac{n^{2}\pi^{2}t_{1}}{L^{2}}}\right)\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}$
$\leq 16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\left|1\wedge\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}\right|^{2}\frac{L^{2}}{n^{2}\pi^{2}}$
$\leq\frac{16e^{2|\alpha|T}}{L}\sum_{n=1}^{\infty}\left(\frac{L^{4}}{n^{4}\pi^{4}}\wedge|t_{2}-t_{1}|^{2}\right).$
Moreover, we obtain
$\frac{16e^{2|\alpha|T}}{L}\sum_{n=1}^{\infty}\left(\frac{L^{4}}{n^{4}\pi^{4}}\wedge|t_{2}-t_{1}|^{2}\right)$
$\leq\frac{16e^{2|\alpha|T}}{L}\left(\sum_{n\leq|t_{2}-t_{1}|^{-1/2}L/\pi}|t_{2}-t_{1}|^{2}+\frac{L^{4}}{\pi^{4}}\sum_{n>|t_{2}-t_{1}|^{-1/2}L/\pi}\frac{1}{n^{4}}\right)$
$\leq\frac{16e^{2|\alpha|T}}{L}\left(\frac{L}{\pi}|t_{2}-t_{1}|^{3/2}+\frac{L^{4}}{\pi^{4}}\int_{|t_{2}-t_{1}|^{-1/2}L/(2\pi)}^{\infty}\frac{1}{y^{4}}\ \mathrm{d}y\right)$
$=\frac{176}{3\pi}e^{2|\alpha|T}|t_{2}-t_{1}|^{3/2}.$
Then, we have
$\int_{0}^{t_{1}}\int_{0}^{L}(\mathcal{I}_{0}(t_{2}-t_{1}+s,y)-\mathcal{I}_{0}(s,y))^{2}\ \mathrm{d}y\mathrm{d}s$
$\leq\frac{16}{6}L|\alpha|^{2}Te^{4|\alpha|T}|t_{2}-t_{1}|^{2}+\frac{176}{3\pi}e^{2|\alpha|T}|t_{2}-t_{1}|^{3/2}$
$\leq\left(\frac{16}{6}L|\alpha|^{2}T^{2}e^{4|\alpha|T}+\frac{176}{3\pi}T^{1/2}e^{2|\alpha|T}\right)|t_{2}-t_{1}|.$
Hence, from (14) and Lemma 3.1, we obtain
$\|F_{L}(t_{2})-F_{L}(t_{1})\|_{k}^{2}$
$\leq\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L^{2}}\left(\int_{t_{1}}^{t_{2}}\int_{0}^{L}\mathcal{I}_{0}^{2}(t_{2}-s,y)\ \mathrm{d}y\mathrm{d}s\right.$
$\quad+\left.\int_{0}^{t_{1}}\int_{0}^{L}(\mathcal{I}_{0}(t_{2}-t_{1}+s,y)-\mathcal{I}_{0}(s,y))^{2}\ \mathrm{d}y\mathrm{d}s\right)$
$\leq\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L^{2}}\left(Le^{2|\alpha|T}|t_{2}-t_{1}|+\left(\frac{16}{6}L|\alpha|^{2}T^{2}e^{4|\alpha|T}+\frac{176}{3\pi}T^{1/2}e^{2|\alpha|T}\right)|t_{2}-t_{1}|\right)$
$\leq\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L}\left(e^{2|\alpha|T}+\frac{16}{6}|\alpha|^{2}T^{2}e^{4|\alpha|T}+\frac{176}{3\pi}T^{1/2}e^{2|\alpha|T}\right)|t_{2}-t_{1}|.$
Therefore, in the case of any boundary conditions, we have (12). This completes the proof. ∎
Proof of Theorem 2.
The tightness of $\{\sqrt{L}F_{L}\}_{L\geq 1}$ in $C([0,T])$ is a consequence of Proposition 5.1 (see [K98]). Then, it remains to prove that the finite-dimensional distributions of the process $(\sqrt{L}F_{L}(t))_{t\in[0,T]}$ converge to those of $(\int_{0}^{t}e^{-\alpha(t-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s})_{t\in[0,T]}$ as $L\to\infty$.
Let us fix $T>0$ and $m\geq 1$ points $t_{1},\dots,t_{m}\in(0,T]$. Proposition 3.1 ensures that, for every $i,j=1,\dots,m$,
$\operatorname{Cov}\left[F_{L}(t_{i}),F_{L}(t_{j})\right]\sim\frac{1}{L}\int_{0}^{t_{i}\wedge t_{j}}e^{-\alpha(t_{i}+t_{j}-2s)}f_{\sigma}(s)\ \mathrm{d}s$ (15)
as $L\to\infty$. We define the following quantities:
$F_{i}:=\frac{F_{L}(t_{i})}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}}$
and
$C_{i,j}:=\operatorname{Cov}\left[F_{i},F_{j}\right]$
for $i,j=1,\dots,m$. We will write $F:=(F_{1},\dots,F_{m})$, and let $G=(G_{1},\dots,G_{m})$ denote a centered Gaussian random vector with covariance matrix $C=(C_{i,j})_{1\leq i,j\leq m}$.
Recall from (9) the random fields $v_{L,t_{1}},\dots,v_{L,t_{m}}$, and define rescaled random fields $V_{1},\dots,V_{m}$ as follows:
$V_{i}:=\frac{v_{L,t_{i}}}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}}$
for $i=1,\dots,m$. According to (8), $F_{i}=\delta(V_{i})$ for all $i=1,\dots,m$. And, by duality, $E[\langle DF_{i},V_{j}\rangle_{\mathcal{H}}]=E[F_{i}\delta(V_{j})]=C_{i,j}$ for all $i,j=1,\dots,m$. Then, Proposition 2.2 implies that
$\left|E[h(F)]-E[h(G)]\right|\leq\frac{1}{2}\|h^{\prime\prime}\|_{\infty}\sqrt{\sum_{i,j=1}^{m}\operatorname{Var}\left(\langle DF_{i},V_{j}\rangle_{\mathcal{H}}\right)}$
for all $h\in C^{2}(\mathbb{R}^{m})$. From Proposition 4.1,
$\operatorname{Var}\left(\langle DF_{i},V_{j}\rangle_{\mathcal{H}}\right)$ $=\frac{\operatorname{Var}\left(\langle DF_{L}(t_{i}),v_{L,t_{j}}\rangle_{\mathcal{H}}\right)}{\operatorname{Var}(F_{L}(t_{i}))\operatorname{Var}(F_{L}(t_{j}))}$
$\leq\frac{A_{T}}{L^{3}\operatorname{Var}(F_{L}(t_{i}))\operatorname{Var}(F_{L}(t_{j}))},$
which together with (15) implies that
$\lim_{L\to\infty}\left|E[h(F)]-E[h(G)]\right|=0$
for all $h\in C^{2}(\mathbb{R}^{m})$.
On the other hand, owing to (15), as $L\to\infty$,
$C_{i,j}=\frac{\operatorname{Cov}(F_{L}(t_{i}),F_{L}(t_{j}))}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}\sqrt{\operatorname{Var}(F_{L}(t_{j}))}}\to\frac{\int_{0}^{t_{i}\wedge t_{j}}e^{-\alpha(t_{i}+t_{j}-2s)}f_{\sigma}(s)\ \mathrm{d}s}{\sqrt{\int_{0}^{t_{i}}e^{-2\alpha(t_{i}-s)}f_{\sigma}(s)\ \mathrm{d}s}\sqrt{\int_{0}^{t_{j}}e^{-2\alpha(t_{j}-s)}f_{\sigma}(s)\ \mathrm{d}s}},$
which yields that the Gaussian random vector $G$ converges weakly to
$\left(\frac{\int_{0}^{t_{1}}e^{-\alpha(t_{1}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{1}}e^{-2\alpha(t_{1}-s)}f_{\sigma}(s)\ \mathrm{d}s}},\dots,\frac{\int_{0}^{t_{m}}e^{-\alpha(t_{m}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{m}}e^{-2\alpha(t_{m}-s)}f_{\sigma}(s)\ \mathrm{d}s}}\right)$ (16)
as $L\to\infty$.
Therefore, it follows that $F$ converges weakly to the random vector in (16) as $L\to\infty$. Then, from (15), it follows that
$\sqrt{L}\left(\frac{F_{L}(t_{1})}{\sqrt{\int_{0}^{t_{1}}e^{-2\alpha(t_{1}-s)}f_{\sigma}(s)\ \mathrm{d}s}},\dots,\frac{F_{L}(t_{m})}{\sqrt{\int_{0}^{t_{m}}e^{-2\alpha(t_{m}-s)}f_{\sigma}(s)\ \mathrm{d}s}}\right)$
converges to the random vector in (16) as $L\to\infty$. This completes the proof. ∎
6 On the limit in Assumption 1
If $\sigma(u)=\sigma_{1}u+\sigma_{0}$ for some constants $\sigma_{1}$ and $\sigma_{0}$, we can calculate the limit $f_{\sigma}(t)$ in Assumption 1. Our calculation is based on the Wiener chaos decomposition of $u(t,x)$. Let
$u_{0}(t,x)=\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y$
for every $(0,T]\times[0,L]$ and $u_{0}(0,x)=u_{0}(x)=1$ for all $x\in[0,L]$. Let
$\mathcal{I}_{0}(t,x):=u_{0}(t,x)$
and
$\mathcal{I}_{k}(t,x):=$ $\sigma_{1}^{k-1}\int_{0}^{t}\int_{0}^{L}\dots\int_{0}^{r_{k-1}}\int_{0}^{L}G_{t-r_{1}}(x,z_{1})\dots G_{r_{k-1}-r_{k}}(z_{k-1},z_{k})$
×$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})\ W(\mathrm{d}r_{k},\mathrm{d}z_{k})\dots W(\mathrm{d}r_{1},\mathrm{d}z_{1})$
for all $k\geq 1$. We define the Picard iteration $\{u_{n}(t,x)\}_{n=0}^{\infty}$ for $u(t,x)$. Define iteratively, for every $n\geq 0$,
$u_{n+1}(t,x):=u_{0}(t,x)+\int_{0}^{t}\int_{0}^{L}G_{t-r}(x,z)\sigma(u_{n}(r,z))\ W(\mathrm{d}r,\mathrm{d}z).$ (17)
Since we only consider the case $\sigma(u)=\sigma_{1}u+\sigma_{0}$ through this section, we have
$u_{n+1}(t,x)$
$=u_{0}(t,x)+\int_{0}^{t}\int_{0}^{L}G_{t-r_{1}}(x,z_{1})(\sigma_{1}u_{n}(r_{1},z_{1})+\sigma_{0})\ W(\mathrm{d}r_{1},\mathrm{d}z_{1}).$
Then, by using mathematical induction, we obtain the Wiener chaos decomposition
$u_{n+1}(t,x)=\sum_{k=0}^{n}\mathcal{I}_{k}(t,x).$ (18)
Therefore, $u_{n}\to u$ in $L^{p}$ ($p\geq 2$) (see [W86]) and applying (18), we have
$u(t,x)=\sum_{k=0}^{\infty}\mathcal{I}_{k}(t,x).$ (19)
Moreover, by multiple Itô’s isometry,
$\|u(t,x)\|_{2}^{2}=\sum_{k=0}^{\infty}\|\mathcal{I}_{k}(t,x)\|_{2}^{2}$ (20)
where
$\|\mathcal{I}_{k}(t,x)\|_{2}^{2}=$ $\sigma_{1}^{2(k-1)}\int_{0}^{t}\int_{0}^{L}\dots\int_{0}^{r_{k-1}}\int_{0}^{L}G_{t-r_{1}}^{2}(x,z_{1})\dots G_{r_{k-1}-r_{k}}^{2}(z_{k-1},z_{k})$
×$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
for all $k\geq 1$.
Using Lemma 3.1, we get the following Propositions.
Proposition 6.1.
Fix $T>0$. In the case of Neumann/Dirichlet boundary conditions, for every $k\geq 1$
$\sup_{L\geq 1}\sup_{(t,x)\in[0,T]\times[0,L]}\|\mathcal{I}_{k}(t,x)\|\leq\sigma_{1}^{2(k-1)}(|\sigma_{1}|e^{|\alpha|T}+|\sigma_{0}|)^{2}\frac{K_{T}^{2k}(4\beta)^{-k/2}T^{k/2}}{\Gamma((k+2)/2)},$ (21)
where the constant $K_{T}$ is defined in (29). Moreover, for every $t>0$ and $k\geq 1$, there exists $f_{k}(t)$ such that
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}\|\mathcal{I}_{k}(t,x)\|_{2}^{2}\ \mathrm{d}x=f_{k}(t).$
Proof.
$\|\mathcal{I}_{k}(t,x)\|_{2}^{2}$
$\leq\sigma_{1}^{2(k-1)}K_{T}^{2k}(|\sigma_{1}|e^{|\alpha|T}+|\sigma_{0}|)^{2}$
×$\quad\times\int_{0}^{t}\int_{\mathbb{R}}\dots\int_{0}^{r_{k-1}}\int_{\mathbb{R}}p_{\beta(t-r_{1})}^{2}(x-z_{1})\dots p_{\beta(r_{k-1}-r_{k})}^{2}(z_{k-1}-z_{k})\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
$=\sigma_{1}^{2(k-1)}K_{T}^{2k}(|\sigma_{1}|e^{|\alpha|T}+|\sigma_{0}|)^{2}\left(\frac{t}{4\pi\beta}\right)^{k/2}$
×$\quad\times\int_{0<r_{k}<\dots<r_{1}<1}\sqrt{\frac{1}{(1-r_{1})\times\dots\times(r_{k-1}-r_{k})}}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
$=\sigma_{1}^{2(k-1)}K_{T}^{2k}(|\sigma_{1}|e^{|\alpha|T}+|\sigma_{0}|)^{2}\left(\frac{t}{4\pi\beta}\right)^{k/2}\frac{\Gamma(1/2)^{k}}{\Gamma((k+2)/2)}$
$\leq\sigma_{1}^{2(k-1)}(|\sigma_{1}|e^{|\alpha|T}+|\sigma_{0}|)^{2}\frac{K_{T}^{2k}(4\beta)^{-k/2}T^{k/2}}{\Gamma((k+2)/2)}$
where the first equality follows from semigroup property of Gaussian heat kernel $p_{t}(z)$ and change of variables, and the second one holds by the following identity
$\int_{0<r_{k}<\dots<r_{1}<1}\sqrt{\frac{1}{(1-r_{1})\times\dots\times(r_{k-1}-r_{k})}}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}=\frac{\Gamma(1/2)^{k}}{\Gamma((k+2)/2)},$ (22)
see [OLBC10]*5.14.1. Then, we have (21).
Using the same arguments as [P22]*Proposition 3.4., we obtain
$\frac{1}{L}\int_{0}^{L}\|\mathcal{I}_{k}(t,x)\|_{2}^{2}\ \mathrm{d}x$
$=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}e^{-2\alpha(t-r_{1})}p_{2\beta(t-r_{1})}(0)\dots e^{-2\alpha(r_{k-1}-r_{k})}p_{2\beta(r_{k-1}-r_{k})}(0)$
×$\quad\times\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
$\quad+o(L).$
Therefore, using dominated convergence theorem, (21) and Lemma 3.1, we have
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}\|\mathcal{I}_{k}(t,x)\|_{2}^{2}\ \mathrm{d}x$ (23)
$=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}e^{-2\alpha(t-r_{1})}p_{2\beta(t-r_{1})}(0)\dots e^{-2\alpha(r_{k-1}-r_{k})}p_{2\beta(r_{k-1}-r_{k})}(0)$
×$\quad\times\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
$=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}e^{-2\alpha(t-r_{1})}p_{2\beta(t-r_{1})}(0)\dots e^{-2\alpha(r_{k-1}-r_{k})}p_{2\beta(r_{k-1}-r_{k})}(0)$
×$\quad\times(\sigma_{1}^{2}e^{-2\alpha r_{k}}+2\sigma_{1}\sigma_{0}e^{-\alpha r_{k}}+\sigma_{0}^{2})\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
$=\sigma_{1}^{2(k-1)}e^{-2\alpha t}\int_{0<r_{k}<\dots<r_{1}<t}p_{2\beta(t-r_{1})}(0)\dots p_{2\beta(r_{k-1}-r_{k})}(0)$
×$\quad\times(\sigma_{1}+e^{\alpha r_{k}}\sigma_{0})^{2}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
$=:f_{k}(t).$
Proposition 6.2.
In the case of periodic boundary conditions, for all $k\geq 1$ and $t>0$, $\|\mathcal{I}_{k}(t,x)\|_{2}^{2}$ does not depend on $x\in[0,L]$ and
$\lim_{L\to\infty}\|\mathcal{I}_{k}(t,x)\|_{2}^{2}=f_{k}(t)$
where $f_{k}(t)$ is the limit in .
Proof.
Since $\mathcal{I}_{0}(t,x)=e^{-\alpha t}$(see Lemma 3.1 and (28) in Lemma A.1), we have
$\|\mathcal{I}_{k}(t,x)\|_{2}^{2}$
$=\sigma_{1}^{2(k-1)}\int_{0}^{t}\int_{0}^{L}\dots\int_{0}^{r_{k-1}}\int_{0}^{L}G_{t-r_{1}}^{2}(x,z_{1})\dots G_{r_{k-1}-r_{k}}^{2}(z_{k-1},z_{k})$
×$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
$=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}G_{2(t-r_{1})}(0,0)\dots G_{2(r_{k-1}-r_{k})}(0,0)(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
where in the second equality follows from semigroup property of $G$. Then, it follows that $\|\mathcal{I}_{k}(t,x)\|_{2}^{2}$ does not depend on $x\in[0,L]$. Moreover, by dominated convergence theorem,
$\lim_{L\to\infty}\|\mathcal{I}_{k}(t,x)\|_{2}^{2}$
$=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}e^{-2\alpha(t-r_{1})}p_{2\beta(t-r_{1})}(0)\dots e^{-2\alpha(r_{k-1}-r_{k})}p_{2\beta(r_{k-1}-r_{k})}(0)$
×$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
$=\sigma_{1}^{2(k-1)}e^{-2\alpha t}\int_{0<r_{k}<\dots<r_{1}<t}p_{2\beta(t-r_{1})}(0)\dots p_{2\beta(r_{k-1}-r_{k})}(0)$
×$\quad\times(\sigma_{1}+\sigma_{0}e^{\alpha r_{k}})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
$=f_{k}(t),$
where $f_{k}$ is the same as in Proposition 6.1. ∎
By Proposition 6.1 and 6.2, we can check the Assumption 1.
Proposition 6.3.
For every $t>0$,
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[\sigma(u(t,x))^{2}]\ \mathrm{d}x$
$=\sigma_{1}^{2}\sum_{k=0}^{\infty}f_{k}(t)+2\sigma_{1}\sigma_{0}e^{-\alpha t}+\sigma_{0}^{2}$
$=:f_{\sigma}(t),$
where $\sigma(u)=\sigma_{1}u+\sigma_{0}$, $f_{0}(t):=e^{-2\alpha t}$ and $f_{k}$ is defined in (23). Moreover,
$f_{\sigma}(t)$ $\leq e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}(f(\sigma_{1}^{4}t/\beta)-1)$
$\quad+(\sigma_{1}e^{-\alpha t}+\sigma_{0})^{2},$
where
$f(t):=2e^{t/4}\int_{-\infty}^{\sqrt{t/2}}\frac{1}{\sqrt{2\pi}}e^{-y^{2}/2}\ \mathrm{d}y.$ (24)
In particular, $f_{\sigma}\in L^{1}([0,T])$ for all $T>0$.
Proof.
We first consider the Neumann/Dirichlet case. By proposition (20) and Tonelli’s theorem,
$\frac{1}{L}\int_{0}^{L}E[u(t,x)^{2}]\ \mathrm{d}x=\sum_{k=0}^{\infty}\frac{1}{L}\int_{0}^{L}\|\mathcal{I}_{k}(t,x)\|_{2}^{2}\ \mathrm{d}x.$
Since the series $\sum_{k=1}^{\infty}\sigma_{1}^{2(k-1)}(|\sigma_{1}|e^{|\alpha|T}+|\sigma_{0}|)^{2}\frac{K_{T}^{2k}(4\beta)^{-k/2}T^{k/2}}{\Gamma((k+2)/2)}$ converges, by Proposition 6.1, dominated convergence theorem and Lemma 3.1, we have
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[u(t,x)^{2}]\ \mathrm{d}x$
$=\sum_{k=0}^{\infty}\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}\|\mathcal{I}_{k}(t,x)\|_{2}^{2}\ \mathrm{d}x$
$=e^{-2\alpha t}+\sum_{k=1}^{\infty}f_{k}(t)$
$=\sum_{k=0}^{\infty}f_{k}(t),$
where $f_{0}(t):=e^{-2\alpha t}$.
Similarly, in the case of periodic boundary conditions, Proposition 6.2 and (20) imply that $E[u(t,x)^{2}]$ does not depend on $x\in[0,L]$. Hence, we have
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[u(t,x)^{2}]\ \mathrm{d}x$
$=\lim_{L\to\infty}E[u(t,0)^{2}]$
$=\sum_{k=0}^{\infty}\lim_{L\to\infty}\|\mathcal{I}_{k}(t,0)\|_{2}^{2}$
$=e^{-2\alpha t}+\sum_{k=1}^{\infty}f_{k}(t)$
$=\sum_{k=0}^{\infty}f_{k}(t),$
where in the second equality, we apply dominated convergence theorem in order to exchange the limit and the sum.
Therefore, in the case of any boundary conditions, we obtain
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[u(t,x)^{2}]\ \mathrm{d}x=\sum_{k=0}^{\infty}f_{k}(t).$
Hence, we have
$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[\sigma(u(t,x))^{2}]\ \mathrm{d}x$
$=\sigma_{1}^{2}\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[u(t,x)^{2}]\ \mathrm{d}x$
$\quad+2\sigma_{1}\sigma_{0}\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[u(t,x)]\ \mathrm{d}x+\sigma_{0}^{2}$
$=\sigma_{1}^{2}\sum_{k=0}^{\infty}f_{k}(t)+2\sigma_{1}\sigma_{0}e^{-\alpha t}+\sigma_{0}^{2},$
where the second equality holds by $E[u(t,x)]=\mathcal{I}_{0}(t,x)$ and Lemma 3.1.
From (22), (23) and change of variables, for all $k\geq 1$, we have
$\sigma_{1}^{2}f_{k}(t)$ $\leq\sigma_{1}^{2k}e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}$
×$\quad\times\int_{0<r_{k}<\dots<r_{1}<t}p_{2\beta(t-r_{1})}(0)\dots p_{2\beta(r_{k-1}-r_{k})}(0)\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
$=e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}\frac{(\sigma_{1}^{4}t/(4\beta))^{k/2}}{\Gamma((k+2)/2)}.$
By using the following identity(see [C13]*Lemma 2.3.4)
$\sum_{n=1}^{\infty}\frac{\lambda^{n-1}}{\Gamma((n+1)/2)}=2e^{\lambda^{2}}\int_{-\infty}^{\sqrt{2}\lambda}\frac{1}{\sqrt{2\pi}}e^{-y^{2}/2}\ \mathrm{d}y,\ \mathrm{for~{}all~{}}\lambda\geq 0$
with $\lambda=\sqrt{\sigma_{1}^{4}t/(4\beta)}$, we obtain
$\sum_{k=0}^{\infty}\frac{(\sigma_{1}^{4}t/(4\beta))^{k/2}}{\Gamma((k+2)/2)}$
$=\sum_{n=1}^{\infty}\frac{(\sigma_{1}^{4}t/(4\beta))^{(n-1)/2}}{\Gamma((n+1)/2)}$
$=f(\sigma_{1}^{4}t/\beta),$
where $f$ is defined in (24). Hence, we have
$\sigma_{1}^{2}\sum_{k=0}^{\infty}f_{k}(t)$
$=\sigma_{1}^{2}e^{-2\alpha t}+\sum_{k=1}^{\infty}\sigma_{1}^{2}f_{k}(t)$
$\leq\sigma_{1}^{2}e^{-2\alpha t}+e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}\sum_{k=1}^{\infty}\frac{(\sigma_{1}^{4}t/(4\beta))^{k/2}}{\Gamma((k+2)/2)}$
$=\sigma_{1}^{2}e^{-2\alpha t}+e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}\sum_{k=0}^{\infty}\frac{(\sigma_{1}^{4}t/(4\beta))^{k/2}}{\Gamma((k+2)/2)}$
$\quad-e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}$
$=\sigma_{1}^{2}e^{-2\alpha t}+e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}(f(\sigma_{1}^{4}t/\beta)-1).$
Therefore, we get
$f_{\sigma}(t)$ $=\sigma_{1}^{2}\sum_{k=0}^{\infty}f_{k}(t)+2\sigma_{1}\sigma_{0}e^{-\alpha t}+\sigma_{0}^{2}$
$\leq e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}(f(\sigma_{1}^{4}t/\beta)-1)$
$\quad+(\sigma_{1}e^{-\alpha t}+\sigma_{0})^{2}.$
Then, $f_{\sigma}\in L^{1}([0,T])$ for all $T>0$. ∎
Appendix A Properties of Green’s function
Denote the Gaussian heat kernel on $\mathbb{R}$ as
$p_{t}(z)=\frac{1}{\sqrt{2\pi t}}e^{-\frac{z^{2}}{2t}},\quad t>0,z\in\mathbb{R}.$ (25)
Let $G$ (resp. $g$) be the Green’s function for the cable equation(resp. heat equation) with Neumann/Dirichlet/periodic boundary conditions. Note that
$G_{t}(x,y)=e^{-\alpha t}g_{\beta t}(x,y).$ (26)
Hence, the properties of $G$ follow directly from those of $g$. For all $t>0$ and $x,y\in[0,L]$, in the case of Neumann boundary conditions,
$G_{t}(x,y)=e^{-\alpha t}\sum_{n\in\mathbb{Z}}\left(p_{\beta t}(x-y+2nL)+p_{\beta t}(x+y+2nL)\right),$
or equivalently,
$G_{t}(x,y)=\frac{e^{-\alpha t}}{\sqrt{L}}+\frac{2e^{-\alpha t}}{L}\sum_{n=1}^{\infty}\cos\left(\frac{n\pi x}{L}\right)\cos\left(\frac{n\pi y}{L}\right)e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}};$
and in the case of Dirichlet boundary conditions,
$G_{t}(x,y)=e^{-\alpha t}\sum_{n\in\mathbb{Z}}\left(p_{\beta t}(x-y+2nL)-p_{\beta t}(x+y+2nL)\right),$
or equivalently,
$G_{t}(x,y)=\frac{2e^{-\alpha t}}{L}\sum_{n=1}^{\infty}\sin\left(\frac{n\pi x}{L}\right)\sin\left(\frac{n\pi y}{L}\right)e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}};$ (27)
and in the case of periodic boundary conditions,
$G_{t}(x,y)=e^{-\alpha t}\sum_{nin\mathbb{Z}}p_{\beta t}(x-y+nL).$
Moreover, by using the properties of $g$ [P22]*Lemma A.1., we get some useful properties of $G$.
Lemma A.1.
  • (1)
    Symmetry. $G_{t}(x,y)=G_{t}(y,x)$ for all $t>0$, $x,y\in[0,L]$.
  • (2)
    In the case of Neumann and periodic boundary conditions, for all $t>0$ and $x\in[0,L]$,
    $\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y=e^{-\alpha t}.$ (28)
  • (3)
    Semigroup property. For all $t,s>0$ and $x,y\in[0,L]$,
    $\int_{0}^{L}G_{t}(x,z)G_{s}(z,y)\ \mathrm{d}z=G_{t+s}(x,y).$
  • (4)
    In the case of Neumann/Dirichlet boundary conditions, for every $t>0$ and $x,y\in[0,L]$,
    $G_{t}(x,y)\leq e^{-\alpha t}p_{\beta t}(x-y)\left(4+\frac{4}{1-e^{-\frac{L^{2}}{\beta t}}}\right),$
    and as a consequence, for all $t\in(0,T]$, $L\geq 1$ and $x,y\in[0,L]$,
    $G_{t}(x,y)\leq K_{T}p_{\beta t}(x-y),$
    where
    $K_{T}=e^{|\alpha|T}\left(4+\frac{4}{1-e^{-\frac{1}{\beta T}}}\right).$ (29)

### v2

@@
 Abstract
-In this paper, we consider one-dimensional nonlinear stochastic cable equations driven by a multiplicative space-time white noise.
-Using the Malliavin–Stein’s method, we prove the Gaussian fluctuation for the spatial average of the solution under the total variation distance with mild conditions.
-We also establish functional central limit theorem.
+We study one-dimensional nonlinear stochastic cable equations driven by a multiplicative space-time white noise.
+Using the Malliavin–Stein method, we prove a central limit theorem for the spatial average of the solution.
+The convergence is established in the total variation distance with mild conditions.
+We also establish a functional central limit theorem with a technical assumption.
+Furthermore, we show that this assumption holds in a special case.
 Keywords: Stochastic cable equation, Central limit theorem, Malliain calculus, Stein’s method 0002020 Mathematics Subject Classification: Primary 60H15; Secondary 60F05, 60H07.
 1 Introduction and main results
-We consider the nonlinear stochastic cable equation
+In this paper, we consider the nonlinear stochastic cable equation
 $\frac{\partial u}{\partial t}=\frac{\beta}{2}\frac{\partial^{2}u}{\partial x^{2}}-\alpha u+\sigma(u)\dot{W}$ (1)
-on $[0,T]\times[0,L]$ for some constants $\alpha\in\mathbb{R}$ and $\beta>0$, where $T>0$ is fixed, $\dot{W}$ is a space-time white noise on $[0,T]\times[0,L]$, with initial condition $u_{0}(x)=1$, Neumann, Dirichlet, or periodic boundary conditions.
-We assume that the coefficient $\sigma\in C^{1}(\mathbb{R})$ satisfies global Lipschitz and linear growth conditions.
-Following [W86], this equation has a unique mild solution, which is adapted to the filtration generated by $W$, such that $E[u(t,x)^{2}]<\infty$ and it satisfies the evolution equation
+on $0,T]\times[0,L]$ for some constants $\alpha\in\mathbb{R}$ and $\beta>0$, where $T>0$ is fixed, $\dot{W}$ is a space-time white noise on $0,T]\times[0,L]$, with initial condition $u_{0}(x)=1$, Neumann, Dirichlet, or periodic boundary conditions.
+We assume the coefficient $\sigma$ is global Lipschitz.
+According to Walsh [W86], the above equation admits a unique mild solution, which is adapted to the filtration generated by $W$ and satisfies the condition $E[u(t,x)^{2}]<\infty$.
+The mild solution satisfies the following equation
 $u(t,x)=\int_{0}^{L}u_{0}(y)G_{t}(x,y)\ \mathrm{d}y+\int_{0}^{t}\int_{0}^{L}G_{t-s}(x,y)\sigma(u(s,y))\ W(\mathrm{d}s,\mathrm{d}y)$ (2)
 where in the right hand side the stochastic integral is in the sense of Itô–Walsh, and $G$ is the Green’s function for the cable equation (5).
-We assume the following technical condition, which is required for the proof of the central limit theorem.
-Assumption 1.
-There exists some positive valued measurable function $f_{\sigma}$ such that for any $T>0$, $f_{\sigma}\in L^{1}([0,T])$ and for all $t\in[0,T]$,
-$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[\sigma(u(t,x))^{2}]\ \mathrm{d}x=f_{\sigma}(t).$
-In this paper, we are interested in the asymptotic behavior as $L$ tends to infinity of the quantity
-$F_{L}(t):=\frac{1}{L}\int_{0}^{L}\left\{u(t,x)-E[u(t,x)]\right\}\ \mathrm{d}x,$ (3)
-where $L\geq 1$, $u(t,x)$ is the solution to (1).
-Recall that the total variation distance between two random variables $X$ and $Y$ is defined as
-$d_{\text{TV}}(X,Y)=\sup_{B\in\mathcal{B}(\mathbb{R})}\left|P(X\in B)-P(Y\in B)\right|$
-where $\mathcal{B}(\mathbb{R})$ denotes the Borel $\sigma$-algebra with respect to the Euclidean topology on $\mathbb{R}$.
-We also write $d_{\text{TV}}(F,\mathcal{N}(0,1))$ for the total variation distance between the law of $F$ and the standard normal law.
+We study the large $L$ asymptotics of the spatial average $F_{L}(t)$ of the solution, given by
+$F_{L}(t):=\frac{1}{L}\int_{0}^{L}\left\{u(t,x)-E[u(t,x)]\right\}\ \mathrm{d}x.$ (3)
+For two random variables $X$ and $Y$, the total variation distance is defined as
+$d_{\text{TV}}(X,Y)=\sup_{B\in\mathcal{B}(\mathbb{R})}|P(X\in B)-P(Y\in B)|,$
+where the supremum is taken over all sets $B$ in the Borel $\sigma$-algebra $\mathcal{B}(\mathbb{R})$.
+We denote by $d_{\text{TV}}(F,\mathcal{N}(0,1))$ the total variation distance between the law of a random variable $F$ and the standard normal distribution.
 We are now ready to state the first main result of this paper.
 Theorem 1.
-Suppose that $\operatorname{Var}(F_{L}(t))>0$. Then, for every $t>0$ there exists a real number $c=c(t)>0$ such that for all $L\geq 1$,
+Suppose that $\sigma(1)\neq 0$. Then, for every $t>0$ there exists a real number $c=c(t)>0$ such that for all $L\geq 1$,
 $d_{\mathrm{TV}}\left(\frac{F_{L}(t)}{\sqrt{\operatorname{Var}(F_{L}(t))}},\mathcal{N}(0,1)\right)\leq\frac{c}{\sqrt{L}}.$ (4)
-We also prove the following functional central limit theorem.
+REMARK $1.$
+Condition $\sigma(1)\neq 0$ guarantees that $\operatorname{Var}(F_{L}(t))>0$. This follows from (11), (13), (14), (35), (36) and (37).
+We assume the following technical condition, which is required for the proof of the functional central limit theorem.
+Assumption 1.
+There exists some nonnegative valued measurable function $f_{\sigma}$ such that for any $T>0$, $f_{\sigma}\in L^{1}([0,T])$ and for all $t\in[0,T]$,
+$\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[\sigma(u(t,x))^{2}]\ \mathrm{d}x=f_{\sigma}(t).$
+We state the following functional central limit theorem.
 Theorem 2.
-Fix $T>0$. Then, as $L\to\infty$,
+Fix $T>0$. Suppose that Assumption 1 holds. Then, as $L\to\infty$,
 $\left(\sqrt{L}F_{L}(t)\right)_{t\in[0,T]}\to\left(\int_{0}^{t}e^{-\alpha(t-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}\right)_{t\in[0,T]}$
 where $f_{\sigma}(t)$ is the limit in Assumption 1 and $W=\{W_{s}\}_{s\in[0,T]}$ denotes a standard one-dimensional Brownian motion, and the convergence is in law on the space of continuous functions $C([0,T])$.
-This kind of asymptotic behavior of spatial averages for stochastic partial differential equations has been studied extensively in recent years.
-This line of research was initiated by Huang, Nualart, and Viitasaari in [HNV20], where they established quantitative central limit theorem and its functional version for one-dimensional nonlinear stochastic heat equations driven by space-time white noise using Malliavin–Stein’s method (cf.
-[NP12]).
-After this result, similar central limit theorems for stochastic partial differential equations have been established under various settings.
-On the other hand, many of these studies deal with SPDEs whose solutions exhibit spatial stationarity.
-In our case, as well as in [P22], we consider SPDEs on the interval $[0,L]$, and therefore the solutions are not stationary in space, which requires technical conditions such as Assumption 1 for the proof of the central limit theorem.
-Moreover, the stochastic cable equation studied in this paper can be seen as a generalization of the stochastic heat equation considered in [P22].
-For these reasons, our study can be regarded as a generalization of [P22].
+In recent years, considerable attention has been paid to the asymptotic behavior of spatial averages for solutions to stochastic partial differential equations.
+This area of inquiry was first explored by Huang, Nualart, and Viitasaari in [HNV20].
+Utilizing the Malliavin–Stein method (see [NP12]), they proved a central limit theorem, along with its functional counterpart, for the one-dimensional nonlinear stochastic heat equation with space-time white noise.
+Following their seminal work, analogous central limit theorems for various types of stochastic partial differential equations have been established.
+For example, Huang, Nualart, Viitasaari, and Zheng studied the $d$-dimensional stochastic heat equation with colored noise [HNVZ20].
+There has been much research on the stochastic heat equation under different settings; see, for example, [ANTV22, BY23, CKNP23, KNP21, KY22, NXZ22, NZ20, P22].
+Similar results are also known for the stochastic wave equation; see, for instance, [DNZ20] for the one-dimensional case, [GNZ21] for the two-dimensional case, and [E24, E25] for the three and higher dimensions.
+For other types of SPDEs, Liu and Shen studied a central limit theorem for the solution to an SPDE related to a pseudo-differential operator that generates a stable-like process [LS23].
+The previous study most relevant to our setting is that of Pu [P22].
+Pu considered the one-dimensional stochastic heat equation with boundary conditions on a bounded interval $[0,L]$ and analyzed the solution using its Wiener chaos decomposition.
+In particular, we remark that Assumption 1 is necessary for the proof of the functional central limit theorem when considering SPDEs on a bounded domain, as in the present work and that of Pu [P22].
 Assumption 1 is verified in Section 6 for the case $\sigma(u)=\sigma_{1}u+\sigma_{0}$, where $\sigma_{1}$ and $\sigma_{0}$ are constants.
@@
 In [HNV20] and related works, since the solution to the SPDE under consideration is spatially stationary, the limit in Assumption 1 reduces to $E[\sigma(u(t,0))^{2}]$, and thus technical conditions such as those in Assumption 1 are not required.
-The nonlinear stochastic cable equation emerges within the framework of mathematical models of neurons, where it governs the propagation of electrical signals along the cylindrical structure of neurons.
-Neurons, the fundamental units of the nervous system, operate through an intricate interplay of chemical, biological, and electrical processes.
-In a standard mathematical idealization, neurons are represented as long, thin cylinders, analogous to electrical cables, enabling the application of cable theory.
-For such a cylindrical structure spanning from $x=0$ to $x=L$, we denote the electrical potential at position $x$ and time $t$ by $u(t,x)$, focusing on variations along the spatial coordinate $x$.
-This potential $u(t,x)$ is governed by a system of non-linear PDE’s, called the Hodgkin-Huxley equations, but in certain ranges of values of $u$, they are well-approximated by the cable equation:
+The nonlinear stochastic cable equation arises in the mathematical modeling of neurons, where it describes the propagation of electrical signals along their cylindrical structure.
+Neurons, the fundamental components of the nervous system, operate through a sophisticated interplay of chemical, biological, and electrical phenomena.
+In a common mathematical simplification, a neuron is idealized as a long, thin cylinder, similar to an electrical cable.
+For such a cylinder on the interval $[0,L]$, we assume the electrical potential $u(t,x)$ depends only on position $x\in[0,L]$ and time $t$.
+While this potential $u(t,x)$ is accurately governed by the Hodgkin–Huxley equations, for specific ranges of $u$, these equations can be closely approximated by the cable equation:
 $\frac{\partial u}{\partial t}=\frac{\beta}{2}\frac{\partial^{2}u}{\partial x^{2}}-\alpha u,$ (5)
-where $\beta/2>0$ represents the rate of diffusion within the neuron and $\alpha\in\mathbb{R}$ is the rate at which ions leak across the membrance [KX93].
-The surface of the neuron is covered with synapses, through which it receives impulses of current.
-If the current arriving at $(t,x)$ is denoted by $F(t,x)$, the system satisfies the inhomogeneous PDE:
+with $\beta/2>0$ representing the diffusion rate within the neuron and $\alpha\in\mathbb{R}$ being the rate of ion leakage through the membrane [KX93].
+The neuron’s surface receives current impulses through synapses.
+If this incoming current at $(t,x)$ is represented by $F(t,x)$, the system obeys the inhomogeneous PDE:
 $\frac{\partial u}{\partial t}=\frac{\beta}{2}\frac{\partial^{2}u}{\partial x^{2}}-\alpha u+F.$
-Even when the system is at rest, occasional random impulses may occur, implying that $F$ generally contains a stochastic component.
-In particular, Walsh studied the case where $F$ is compound Poisson process or space-time white noise [W81].
-In [W86], the nonlinear stochastic cable equation was studied for the case $F=\sigma(u)\dot{W}$.
+Even in a resting state, occasional random impulses may occur, which implies that $F$ generally contains a stochastic component.
+For instance, Walsh studied the case where $F$ is a compound Poisson process or a space-time white noise [W81].
+The nonlinear stochastic cable equation with $F=\sigma(u)\dot{W}$ was analyzed in [W86].
 2 Preliminaries
-We write $\|Z\|_{k}$ instead of $E[|Z|^{k}]^{1/k}$, for every $Z\in L^{k}(\Omega)$.
+For any $Z\in L^{k}(\Omega)$, we define $\|Z\|_{k}:=E[|Z|^{k}]^{1/k}$.
+Throughout this paper, we assume that $\sigma$ is globally Lipschitz continuous with constant $K_{\sigma}$. This implies the following linear growth condition:
+$|\sigma(y)|\leq M_{\sigma}(1+|y|)$
+where $M_{\sigma}=\max\{K_{\sigma},|\sigma(0)|\}$.
 2.1 Clark–Ocone formula
-Let $\mathcal{H}=L^{2}([0,T]\times\mathbb{R})$. The Gaussian family $\{X(h)\}_{h\in\mathcal{H}}$ formed by the stochastic integrals
+Let $\mathcal{H}$ be the Hilbert space $L^{2}([0,T]\times\mathbb{R})$. The family of stochastic integrals
 $X(h)=\int_{[0,T]\times\mathbb{R}}h(s,x)\ W(\mathrm{d}s,\mathrm{d}x)$
-defines an isonormal Gaussian process on the Hilbert space $\mathcal{H}$.
-In this framework we can develop the Malliavin calculus [N06].
-We denote by $D$ the Malliavin derivative operator.
-Let $\{\mathcal{F}_{s}\}_{s\in[0,T]}$ denote the filtration generated by the space-time white noise $\dot{W}$.
-We recall the following Clark–Ocone formula:
-$F=E[F]+\int_{[0,T]\times\mathbb{R}}E[D_{s,z}F\mid\mathcal{F}_{s}]\ W(\mathrm{d}s,\mathrm{d}z)\mathrm{~{}a.s.},$
-valid for every $\mathcal{F}_{T}$measurable random variable $F$ in the Sobolev space $\mathbb{D}^{1,2}$ in the sense of Malliavin derivative.
-By using Jensen’s inequality for conditional expectations, the above Clark–Ocone formula yields the following Poincaré-type inequality:
-$\left|\operatorname{Cov}[F,G]\right|\leq\int_{0}^{T}\int_{\mathbb{R}}\|D_{s,z}F\|_{2}\|D_{s,z}G\|_{2}\ \mathrm{d}z\mathrm{d}s$ (6)
-for all $\mathcal{F}_{T}$measurable random variables $F,G\in\mathbb{D}^{1,2}$.
-2.2 The Malliavin–Stein’s method
-Let us now introduce some preliminary results of Malliavin–Stein’s method [NP12].
-Stein’s method is a probabilistic technique to allow one to get some bounds for the distance between two probability measures.
-Combining Stein’s method and the Malliavin calculus, we can obtain useful estimate to prove quantitative central limit theorems.
+constitutes an isonormal Gaussian process $\{X(h)\}_{h\in\mathcal{H}}$.
+In this context, we employ the tools of Malliavin calculus (see, e.g., [N06]), and let $D$ denote the associated Malliavin derivative operator.
+Let $\{\mathcal{F}_{s}\}_{s\in[0,T]}$ be the filtration generated by the space-time white noise $\dot{W}$.
+A key result is the Clark–Ocone formula, which provides the following representation for any $\mathcal{F}_{T}$-measurable random variable $F$ in the Sobolev space $\mathbb{D}^{1,2}$:
+$F=E[F]+\int_{[0,T]\times\mathbb{R}}E[D_{s,z}F\mid\mathcal{F}_{s}]\ W(\mathrm{d}s,\mathrm{d}z)\quad\text{a.s.}$
+A well-known consequence of this formula, obtained via Jensen’s inequality for conditional expectations, is the following Poincaré-type inequality:
+$\left|\operatorname{Cov}[F,G]\right|\leq\int_{0}^{T}\int_{\mathbb{R}}\|D_{s,z}F\|_{2}\|D_{s,z}G\|_{2}\ \mathrm{d}z\mathrm{d}s,$ (6)
+which holds for any pair of $\mathcal{F}_{T}$-measurable random variables $F,G\in\mathbb{D}^{1,2}$.
+2.2 The Malliavin–Stein method
+We now recall some key results from the Malliavin–Stein method.
+This method merges the Malliavin calculus with Stein’s method to provide quantitative bounds on the distance between probability measures, proving particularly effective for central limit theorems [NP12].
 Proposition 2.1.
-Let $F=\delta(v)$ for some $v\in\operatorname{Dom}(\delta)$ which is the $L^{2}(\Omega)$-domain of the adjoint of the Malliavin derivative operator.
+Let $F=\delta(v)$ for some $v\in\operatorname{Dom}(\delta)$, where $\operatorname{Dom}(\delta)$ is the $L^{2}(\Omega)$-domain of the adjoint of the Malliavin derivative operator.
 Suppose that $F\in\mathbb{D}^{1,2}$ and $E[F^{2}]=1$.
-Then we have
+Then, the following inequality holds:
 $d_{\mathrm{TV}}(F,\mathcal{N}(0,1))\leq 2\sqrt{\operatorname{Var}(\langle DF,v\rangle_{\mathcal{H}})}.$
-In order to establish the functional CLT in Theorem 2, we also need the following multivariate counterpart of Proposition 2.1.
+The proof of the functional CLT in Theorem 2 requires a multivariate version of Proposition 2.1, which we state below.
 Proposition 2.2.
-Fix $m\geq 2$.
-Let $F=(F_{1},\dots,F_{m})$ be a random vector such that for every $i=1,\dots,m$, $F_{i}=\delta(v_{i})$ for some $v_{i}\in\operatorname{Dom}(\delta)$ and $F_{i}\in\mathbb{D}^{1,2}$.
-Let $Z$ be an $m$-dimensional centered Gaussian random vector with covariance matrix $(C_{i,j})_{1\leq i,j\leq m}$.
-Then, for every $h\in C^{2}(\mathbb{R}^{m})$ that has bounded second partial derivatives,
+Fix an integer $m\geq 2$.
+Let $F=(F_{1},\dots,F_{m})$ be a random vector where each component has the form $F_{i}=\delta(v_{i})$ for some $v_{i}\in\operatorname{Dom}(\delta)$ and $F_{i}\in\mathbb{D}^{1,2}$.
+Let $Z$ denote an $m$-dimensional centered Gaussian random vector with a given covariance matrix $(C_{i,j})_{1\leq i,j\leq m}$.
+Then, for any function $h\in C^{2}(\mathbb{R}^{m})$ with bounded second partial derivatives, we have
 $\left|E[h(F)]-E[h(Z)]\right|\leq\frac{1}{2}\|h^{\prime\prime}\|_{\infty}\sqrt{\sum_{i,j=1}^{m}E\left[\left|C_{i,j}-\langle DF_{i},v_{j}\rangle_{\mathcal{H}}\right|^{2}\right]},$
@@
 $\|h^{\prime\prime}\|_{\infty}:=\max_{1\leq i,j\leq m}\sup_{x\in\mathbb{R}^{m}}\left|\frac{\partial^{2}h(x)}{\partial x_{i}\partial x_{j}}\right|.$
-2.3 Moments and Malliavin derivative of $u(t,x)$
+2.3 Moments and Malliavin derivative of the solution
 Recently, Pu proved that the moments of $u(t,x)$ are uniformly bounded for all $L\geq 1$ in the case where $\alpha=0$ and $\beta=1$ [P22]*Lemma 2.3..
@@
 Lemma 2.1.
-Let $u(t,x)$ be the solution to (1). Then for all $k\geq 2$,
-$\sup_{L\geq 1}\sup_{(t,x)\in[0,T]\times[0,L]}\|u_{n}(t,x)\|_{k}\leq c_{T,k}<\infty.$ (7)
+Fix $T>0$. Then for all $k\geq 2$, there exists $c_{T,k}>0$ such that
+$\sup_{L\geq 1}\sup_{(t,x)\in[0,T]\times[0,L]}\|u(t,x)\|_{k}\leq c_{T,k}<\infty.$ (7)
 Pu also proved that the moments of the Malliavin derivative of $u(t,x)$ satisfy a Gaussian-type upper bound uniformly over $L\geq 1$ in the case where $\alpha=0$ and $\beta=1$ [P22]*Lemma 2.4..
@@
 Fix $T>0$. Then, for all $k\geq 2$, there exists $C_{T,k}>0$ such that for all $L\geq 1$, $(t,x)\in[0,T]\times[0,L]$, and for almost every $(s,y)\in(0,t)\times\mathbb{R}$,
-$\|D_{s,y}u(t,x)\|_{k}\leq\begin{cases}C_{T,k}1_{[0,L]}(y)p_{\beta(t-s)}(x-y)&\mathrm{(Neumann/Dirichlet~{}case)}\\ C_{T,k}1_{[0,L]}(y)G_{t-s}(x-y)&\mathrm{(periodic~{}case)}\end{cases}$
+$\|D_{s,y}u(t,x)\|_{k}\leq\begin{cases}C_{T,k}1_{[0,L]}(y)p_{\beta(t-s)}(x-y)&\mathrm{(Neumann/Dirichlet~case)}\\ C_{T,k}1_{[0,L]}(y)G_{t-s}(x,y)&\mathrm{(periodic~case)}\end{cases}$
 where $p_{t}(z)$ denotes the Gaussian heat kernel, defined in (33). In particular, we have $u(t,x)\in\bigcap_{k\geq 2}\mathbb{D}^{1,k}$ for all $(t,x)\in[0,T]\times[0,L]$.
@@
 $\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}\mathcal{I}_{0}(t_{1},x)\mathcal{I}_{0}(t_{2},x)\ \mathrm{d}x=e^{-\alpha(t_{1}+t_{2})}.$
-The following result provides the asymptotic behavior of the covariance function of the renormalized sequence of processes $F_{L}(t)$ as $L$ tends to infinity.
+The following lemmas are prepared for the proof of Proposition 3.1.
+Lemma 3.2.
+In the case of Neumann or periodic boundary conditions,
+$\lim_{t\to 0}\sup_{L\geq 1}\sup_{x\in[0,L]}|E[\sigma(u(t,x))^{2}]-\sigma(1)^{2}|=0.$
+Proof.
+Fix $T>0$. For every $(t,x)\in[0,T]\times[0,L]$,
+$|E[\sigma(u(t,x))^{2}]-\sigma(1)^{2}|$
+$=|E[(\sigma(u(t,x))-\sigma(1))(\sigma(u(t,x))+\sigma(1))]|$
+$\leq\sqrt{E[|\sigma(u(t,x))-\sigma(1)|^{2}]}\sqrt{E[|\sigma(u(t,x))+\sigma(1)|^{2}]}$
+$\leq\sqrt{E[K_{\sigma}^{2}|u(t,x)-1|^{2}]}\sqrt{E[2\sigma(u(t,x))^{2}+2\sigma(1)^{2}]}$
+$\leq K_{\sigma}\sqrt{E[|u(t,x)-1|^{2}]}\sqrt{4M_{\sigma}^{2}(c_{T,2}^{2}+1)+2\sigma(1)^{2}},$
+where $K_{\sigma}$ is the Lipschitz constant of $\sigma$, and $M_{\sigma}$ is the constant in the linear growth condition. From (2) and Itô’s isometry, we get
+$E[|u(t,x)-1|^{2}]$ (8)
+$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+2\int_{0}^{t}\int_{0}^{L}G_{t-s}(x,y)^{2}E[\sigma^{2}(u(s,y))]\ \mathrm{d}s\mathrm{d}y$
+$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})\int_{0}^{t}\int_{0}^{L}G_{t-s}(x,y)^{2}\ \mathrm{d}s\mathrm{d}y.$
+Since we assume Neumann or periodic boundary conditions, using (38) and semigroup property of $G$, we have
+$E[|u(t,x)-1|^{2}]$
+$\leq 2|e^{-\alpha t}-1|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})\int_{0}^{t}G_{2(t-s)}(x,x)\ \mathrm{d}s$
+In the case of Neumann boundary conditions, using (39), we get
+$\int_{0}^{t}G_{2(t-s)}(x,x)\ \mathrm{d}s$ $\leq K_{2T}e^{2|\alpha|T}\int_{0}^{t}\frac{1}{\sqrt{4\pi\beta(t-s)}}\ \mathrm{d}s$
+$=K_{2T}e^{2|\alpha|T}\sqrt{\frac{t}{\pi\beta}}.$
+Then, we have
+$E[|u(t,x)-1|^{2}]$
+$\leq 2|e^{-\alpha t}-1|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})K_{2T}e^{2|\alpha|T}\sqrt{\frac{t}{\pi\beta}}$
+In the case of periodic boundary conditions, using (37) and the following identity [P22]*A.17
+$\sum_{j\in\mathbb{Z}}p_{r}(j)=\sum_{n\in\mathbb{Z}}e^{-2r\pi^{2}n^{2}}\quad\text{for}~\text{all}~r>0,$
+we have
+$\int_{0}^{t}G_{2(t-s)}(x,x)\ \mathrm{d}s$
+$=\int_{0}^{t}e^{-2\alpha(t-s)}\sum_{j\in\mathbb{Z}}p_{2\beta(t-s)}(jL)\ \mathrm{d}s$
+$\leq\int_{0}^{t}e^{-2\alpha(t-s)}\sum_{j\in\mathbb{Z}}p_{2\beta(t-s)}(j)\ \mathrm{d}s$
+$=\int_{0}^{t}e^{-2\alpha(t-s)}\sum_{n\in\mathbb{Z}}e^{-4\beta(t-s)\pi^{2}n^{2}}\ \mathrm{d}s$
+$\leq e^{2|\alpha|T}\sum_{n\in\mathbb{Z}}\int_{0}^{t}e^{-4\beta(t-s)\pi^{2}n^{2}}\ \mathrm{d}s$
+$=e^{2|\alpha|T}\sum_{n\in\mathbb{Z}}\frac{1-e^{-4\beta t\pi^{2}n^{2}}}{4\beta\pi^{2}n^{2}}.$
+Then,
+$E[|u(t,x)-1|^{2}]$
+$\leq 2|e^{-\alpha t}-1|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})e^{2|\alpha|T}\sum_{n\in\mathbb{Z}}\frac{1-e^{-4\beta t\pi^{2}n^{2}}}{4\beta\pi^{2}n^{2}}.$
+Applying the Lebesgue dominated convergence theorem, we get
+$\lim_{t\to 0}\sum_{n\in\mathbb{Z}}\frac{1-e^{-4\beta t\pi^{2}n^{2}}}{4\beta\pi^{2}n^{2}}=\sum_{n\in\mathbb{Z}}\lim_{t\to 0}\frac{1-e^{-4\beta t\pi^{2}n^{2}}}{4\beta\pi^{2}n^{2}}=0.$
+Therefore, in the case of Neumann or periodic boundary conditions,
+$\lim_{t\to 0}\sup_{L\geq 1}\sup_{x\in[0,L]}|E[\sigma(u(t,x))^{2}]-E[\sigma(1)^{2}]|=0.$
+Lemma 3.3.
+In the case of Dirichlet boundary conditions,
+$\lim_{t\to 0}\sup_{L\geq 1}\left|\frac{1}{L}\int_{0}^{L}E[\sigma(u(t,x))^{2}]\ \mathrm{d}x-\sigma(1)^{2}\right|=0.$
+Proof.
+Fix $T>0$. For every $(t,x)\in[0,T]\times[0,L]$,
+$\left|\frac{1}{L}\int_{0}^{L}E[\sigma(u(t,x))^{2}]\ \mathrm{d}x-\sigma(1)^{2}\right|$
+$=\left|\frac{1}{L}\int_{0}^{L}E[(\sigma(u(t,x))-\sigma(1))(\sigma(u(t,x))+\sigma(1))]\ \mathrm{d}x\right|$
+$\leq\sqrt{\frac{1}{L}\int_{0}^{L}E[|\sigma(u(t,x))-\sigma(1)|^{2}]\ \mathrm{d}x}\sqrt{\frac{1}{L}\int_{0}^{L}E[|\sigma(u(t,x))+\sigma(1)|^{2}]\ \mathrm{d}x}$
+$\leq\sqrt{\frac{1}{L}\int_{0}^{L}E[K_{\sigma}^{2}|u(t,x)-1|^{2}]\ \mathrm{d}x}\sqrt{\frac{1}{L}\int_{0}^{L}E[2\sigma(u(t,x))^{2}+2\sigma(1)^{2}]\ \mathrm{d}x}$
+$\leq K_{\sigma}\sqrt{\frac{1}{L}\int_{0}^{L}E[|u(t,x)-1|^{2}]\ \mathrm{d}x}\sqrt{4M_{\sigma}^{2}(c_{T,2}^{2}+1)+2\sigma(1)^{2}}.$
+From (2) and Itô’s isometry, we get (8). Since we assume Dirichlet boundary conditions, using semigroup property of $G$ and (39), we have
+$E[|u(t,x)-1|^{2}]$
+$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})\int_{0}^{t}G_{2(t-s)}(x,x)\ \mathrm{d}s$
+$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})K_{2T}e^{2|\alpha|T}\int_{0}^{t}\frac{1}{\sqrt{4\pi\beta(t-s)}}\ \mathrm{d}s$
+$\leq 2\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}+4M_{\sigma}^{2}(1+c_{T,2}^{2})K_{2T}e^{2|\alpha|T}\sqrt{\frac{t}{\pi\beta}}.$
+By (36), we get
+$\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y$
+$=2e^{-\alpha t}\sum_{n=1}^{\infty}\sin\left(\frac{n\pi x}{L}\right)\frac{1-(-1)^{n}}{n\pi}e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}}.$
+Then, applying the $L^{2}([0,L])$-orthogonality of the functions $\{x\mapsto\sin(n\pi x/L)\}_{n=1}^{\infty}$,
+$\frac{1}{L}\int_{0}^{L}\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}\ \mathrm{d}x$ (9)
+$=4e^{-2\alpha t}\sum_{n=1}^{\infty}\frac{1}{L}\int_{0}^{L}\sin^{2}\left(\frac{n\pi x}{L}\right)\ \mathrm{d}x\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}$
+$\quad+1-4e^{-\alpha t}\sum_{n=1}^{\infty}\frac{1}{L}\int_{0}^{L}\sin\left(\frac{n\pi x}{L}\right)\ \mathrm{d}x\frac{1-(-1)^{n}}{n\pi}e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}}$
+$=2e^{-2\alpha t}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}+1-4e^{-\alpha t}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}}.$
+Now, for every $L\geq 1$,
+$\left|\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}-\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}\right|$
+$=\left|\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}(e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}-1)\right|$
+$\leq\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}|e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}-1|$
+$\leq\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}|e^{-n^{2}\pi^{2}\beta t}-1|.$
+Applying the dominated convergence theorem,
+$\lim_{t\to 0}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}|e^{-n^{2}\pi^{2}\beta t}-1|=0.$
+Combining the above with the following identity
+$\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}=\sum_{k=1}^{\infty}\frac{4}{(2k-1)^{2}\pi^{2}}=\frac{1}{2}$
+yields
+$\lim_{t\to 0}\sup_{L\geq 1}\left|\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{L^{2}}}-\frac{1}{2}\right|=0.$
+Similarly, we get
+$\lim_{t\to 0}\sup_{L\geq 1}\left|\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n\pi}\right)^{2}e^{-\frac{n^{2}\pi^{2}\beta t}{2L^{2}}}-\frac{1}{2}\right|=0.$
+These together with (9) imply
+$\lim_{t\to 0}\sup_{L\geq 1}\frac{1}{L}\int_{0}^{L}\left|\int_{0}^{L}G_{t}(x,y)\ \mathrm{d}y-1\right|^{2}\ \mathrm{d}x=0.$
+Therefore, we have
+$\lim_{t\to 0}\sup_{L\geq 1}\frac{1}{L}\int_{0}^{L}E[|u(t,x)-1|^{2}]\ \mathrm{d}x=0.$
+This completes the proof.
+The following results provide the asymptotic behavior of the covariance function of the renormalized sequence of processes $F_{L}(t)$ as $L$ tends to infinity.
 Proposition 3.1.
-For every $t_{1},t_{2}>0$,
-$\lim_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]=\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}f_{\sigma}(s)\ \mathrm{d}s,$
-where the function $f_{\sigma}$ is defined in Assumption 1.
+There exists $\delta>0$, for every $t_{1},t_{2}>0$,
+$\liminf_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]\geq\frac{\sigma(1)^{2}}{2}\int_{0}^{t_{1}\wedge t_{2}\wedge\delta}e^{-\alpha(t_{1}+t_{2}-2s)}\ \mathrm{d}s.$ (10)
 Proof.
 Using the mild form in (2) and Itô’s isometry, we have
-$\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$
+$\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$ (11)
 $=\frac{1}{L}\int_{[0,L]^{2}}\operatorname{Cov}\left[u(t_{1},x),u(t_{2},y)\right]\ \mathrm{d}x\mathrm{d}y$
@@
 $\to 0\mathrm{~as~}L\to\infty.$
-Then, by Assumption 1, we have
+Then, we have
+$\liminf_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$ (12)
+$=\liminf_{L\to\infty}\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}\frac{1}{L}\int_{0}^{L}E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s$
+In the case of Neumann or periodic boundary conditions, from Lemma 3.2, there exists $\delta>0$, for every $L\geq 1$ and $(s,z)\in[0,\delta]\times[0,L]$,
+$E[\sigma(u(s,z))^{2}]\geq\frac{\sigma(1)^{2}}{2}.$ (13)
+In the case of Dirichlet boundary conditions, from Lemma 3.3, there exists $\delta>0$, for every $L\geq 1$ and $s\in[0,\delta]$,
+$\frac{1}{L}\int_{0}^{L}E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\geq\frac{\sigma(1)^{2}}{2}.$ (14)
+These together with (12) implies (10). ∎
+Proposition 3.2.
+Suppose that Assumption 1 holds. For every $t_{1},t_{2}>0$,
+$\lim_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]=\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}f_{\sigma}(s)\ \mathrm{d}s,$ (15)
+where the function $f_{\sigma}$ is defined in Assumption 1.
+Proof.
+By the same argument as in the proof of Proposition 3.1, for every $t_{1},t_{2}>0$,
 $\lim_{L\to\infty}\operatorname{Cov}\left[\sqrt{L}F_{L}(t_{1}),\sqrt{L}F_{L}(t_{2})\right]$
 $=\lim_{L\to\infty}\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}\frac{1}{L}\int_{0}^{L}E[\sigma(u(s,z))^{2}]\ \mathrm{d}z\mathrm{d}s$
-$=\int_{0}^{t_{1}\wedge t_{2}}e^{-\alpha(t_{1}+t_{2}-2s)}f_{\sigma}(s)\ \mathrm{d}s.$
+This together with Assumption 1 implies (15). ∎
 4 Proof of Theorem 1
@@
 $v_{L,t}(s,y)$ (17)
-$:=\frac{1}{L}1_{(0,t)}(s)1_{[0,L]}(y)\sigma(u(s,y))\int_{0}^{L}G_{t-s}(x,y)\ \mathrm{d}x$
+$=\frac{1}{L}1_{(0,t)}(s)1_{[0,L]}(y)\sigma(u(s,y))\int_{0}^{L}G_{t-s}(x,y)\ \mathrm{d}x$
 $=\frac{1}{L}1_{(0,t)}(s)1_{[0,L]}(y)\sigma(u(s,y))\mathcal{I}_{0}(t-s,y),$
@@
 $=\frac{1}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\operatorname{Cov}\left[\sigma(u(s_{1},y_{1}))^{2},\sigma(u(s_{2},y_{2}))^{2}\right]$
-×$\quad\times\mathcal{I}_{0}(t-s_{1},y_{1})\mathcal{I}_{0}(\tau-s_{1},y_{1})\mathcal{I}_{0}(t-s_{2},y_{2})\mathcal{I}_{0}(\tau-s_{2},y_{2})\ \mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}.$
+$\quad\times\mathcal{I}_{0}(t-s_{1},y_{1})\mathcal{I}_{0}(\tau-s_{1},y_{1})\mathcal{I}_{0}(t-s_{2},y_{2})\mathcal{I}_{0}(\tau-s_{2},y_{2})\ \mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}.$
 By Lemma 3.1, Hölder’s inequality and Minkowski inequality, we have
@@
 $\leq\frac{4e^{4|\alpha|T}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{\mathbb{R}}\|\sigma(u(s_{1},y_{1}))D_{r,z}\sigma(u(s_{1},y_{1}))\|_{2}$
-×$\quad\times\|\sigma(u(s_{2},y_{2}))D_{r,z}\sigma(u(s_{2},y_{2}))\|_{2}\ \mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
+$\quad\times\|\sigma(u(s_{2},y_{2}))D_{r,z}\sigma(u(s_{2},y_{2}))\|_{2}\ \mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
 $\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{\mathbb{R}}\|D_{r,z}\sigma(u(s_{1},y_{1}))\|_{4}\|D_{r,z}\sigma(u(s_{2},y_{2}))\|_{4}$
-×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
-where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative [N06],
-$D_{r,z}\sigma(u(s,y))=\sigma^{\prime}(u(s,y))D_{r,z}u(s,y).$
-Then, we obtain
+$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
+where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative for a Lipschitz function with constant $K_{\sigma}$ [N06],
+$D_{r,z}\sigma(u(s,y))=G_{\sigma,u(s,y)}D_{r,z}u(s,y),$
+where $G_{\sigma,u(s,y)}$ is a random variable bounded by $K_{\sigma}$. Then, we obtain
 $\Phi_{L,t,\tau}^{(1)}$
 $\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{\mathbb{R}}\|D_{r,z}u(s_{1},y_{1})\|_{4}\|D_{r,z}u(s_{2},y_{2})\|_{4}$
-×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
+$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
 where $K_{\sigma}$ is Lipschitz constant of $\sigma$. In the case of Neumann/Dirichlet boundary conditions, by Lemma 2.2, we have
@@
 $\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{\mathbb{R}}p_{\beta(s_{1}-r)}(y_{1}-z)p_{\beta(s_{2}-r)}(y_{2}-z)$
-×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
+$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
 $=\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}p_{\beta(s_{1}+s_{2}-2r)}(y_{1}-y_{2})$
-×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
+$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2},$
 where the equality holds by the semigroup property of Gaussian heat kernel $p_{t}(z)$. By using the identity
@@
 $\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]}\int_{\mathbb{R}}\int_{0}^{s_{1}\wedge s_{2}}p_{\beta(s_{1}+s_{2}-2r)}(y_{1})$
-×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
+$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
 $=\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}\int_{[0,t\wedge\tau]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\ \mathrm{d}r\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
 $\leq\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{[0,t\wedge\tau]^{2}}\int_{[0,L]^{2}}\int_{0}^{s_{1}\wedge s_{2}}\int_{0}^{L}G_{s_{1}-r}(y_{1},z)G_{s_{2}-r}(y_{2},r)$
-×$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
+$\quad\times\mathrm{d}z\mathrm{d}r\mathrm{d}y_{1}\mathrm{d}y_{2}\mathrm{d}s_{1}\mathrm{d}s_{2}$
 $=\frac{4e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}\int_{[0,t\wedge\tau]^{2}}\int_{0}^{s_{1}\wedge s_{2}}e^{-\alpha(s_{1}+s_{2}-r)}\mathrm{d}r\mathrm{d}s_{1}\mathrm{d}s_{2}$
@@
 $=\frac{1}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\mathcal{I}_{0}(\tau-r_{1},z_{1})\mathcal{I}_{0}(\tau-r_{2},z_{2})\mathcal{I}_{0}^{2}(t-s,y)$
-×$\quad\times E[\sigma(u(r_{1},z_{1}))(D_{r_{1},z_{1}}\sigma(u(s,y)))\sigma(u(r_{2},z_{2}))(D_{r_{2},z_{2}}\sigma(u(s,y)))]\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s.$
+$\quad\times E[\sigma(u(r_{1},z_{1}))(D_{r_{1},z_{1}}\sigma(u(s,y)))\sigma(u(r_{2},z_{2}))(D_{r_{2},z_{2}}\sigma(u(s,y)))]\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s.$
 By Lemma 3.1, Hölder’s inequality and Minkowski inequality,
@@
 $\leq\frac{e^{4|\alpha|T}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|\sigma(u(r_{1},z_{1}))\|_{4}\|D_{r_{1},z_{1}}\sigma(u(s,y))\|_{4}$
-×$\quad\times\|\sigma(u(r_{2},z_{2}))\|_{4}\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
+$\quad\times\|\sigma(u(r_{2},z_{2}))\|_{4}\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
 $\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|D_{r_{1},z_{1}}\sigma(u(s,y))\|_{4}$
-×$\quad\times\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
-where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative [N06], we obtain
+$\quad\times\|D_{r_{2},z_{2}}\sigma(u(s,y))\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
+where $M_{\sigma}$ is the constant in the linear growth condition for $\sigma$. From the chain rule of Malliavin derivative for a Lipschitz function [N06], we obtain
 $\Phi_{L,t,\tau}^{(2)}$
-$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|\sigma^{\prime}(u(s,y))D_{r_{1},z_{1}}u(s,y)\|_{4}$
-×$\quad\times\|\sigma^{\prime}(u(s,y))D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
+$\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|G_{\sigma,u(s,y)}D_{r_{1},z_{1}}u(s,y)\|_{4}$
+$\quad\times\|\sigma^{\prime}(u(s,y))D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
 $\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\|D_{r_{1},z_{1}}u(s,y)\|_{4}$
-×$\quad\times\|D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
-where $K_{\sigma}$ is Lipschitz constant of $\sigma$. In the case of Neumann/Dirichlet boundary conditions, applying Lemma 2.2, we have
+$\quad\times\|D_{r_{2},z_{2}}u(s,y)\|_{4}\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s,$
+where $K_{\sigma}$ is Lipschitz constant of $\sigma$ and $G_{\sigma,u(s,y)}$ is a random variable bounded by $K_{\sigma}$.
+In the case of Neumann/Dirichlet boundary conditions, applying Lemma 2.2, we have
 $\Phi_{L,t,\tau}^{(2)}$
 $\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}$
-×$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
+$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
 $\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}\int_{\mathbb{R}}$
-×$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}y\mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
+$\quad\times p_{\beta(s-r_{1})}(y-z_{1})p_{\beta(s-r_{2})}(y-z_{2})\ \mathrm{d}y\mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
 $=\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}$
-×$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s,$
+$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s,$
 where the equality holds by the semigroup property of Gaussian heat kernel $p_{t}(z)$. By using the identity (18), we obtain
@@
 $\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}\int_{0}^{L}\int_{\mathbb{R}}$
-×$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
+$\quad\times p_{\beta(2s-r_{1}-r_{2})}(z_{1}-z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
 $=\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
@@
 $\leq\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{4}}\int_{0}^{t}\int_{0}^{L}\int_{[0,\tau\wedge s]^{2}}\int_{[0,L]^{2}}$
-×$\quad\times G_{s-r_{1}}(y,z_{1})G_{s-r_{2}}(y,z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
+$\quad\times G_{s-r_{1}}(y,z_{1})G_{s-r_{2}}(y,z_{2})\ \mathrm{d}z_{1}\mathrm{d}z_{2}\mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}y\mathrm{d}s$
 $=\frac{e^{4|\alpha|T}M_{\sigma}^{2}(1+c_{T,4})^{2}K_{\sigma}^{2}C_{T,4}^{2}}{L^{3}}\int_{0}^{t}\int_{[0,\tau\wedge s]^{2}}e^{-\alpha(2s-r_{1}-r_{2})}\ \mathrm{d}r_{1}\mathrm{d}r_{2}\mathrm{d}s$
@@
 Applying Proposition 4.1 with $t=\tau$, we have
-$\sup_{t\in[0,T]}\operatorname{Var}\left(\left\langle DF_{L}(t),v_{L,t}\right\rangle_{\mathcal{H}}\right)\leq\frac{A_{T}}{L^{3}}\text{~{}for~{}all~{}}L\geq 1$
+$\sup_{t\in[0,T]}\operatorname{Var}\left(\left\langle DF_{L}(t),v_{L,t}\right\rangle_{\mathcal{H}}\right)\leq\frac{A_{T}}{L^{3}}$
 for all $L\geq 1$. Then, by using (16) and Proposition 2.1,
@@
 $\leq 2\sqrt{\operatorname{Var}\left(\left\langle\frac{DF_{L}(t)}{\sqrt{\operatorname{Var}(F_{L}(t))}},\frac{v_{L,t}}{\sqrt{\operatorname{Var}(F_{L}(t))}}\right\rangle_{\mathcal{H}}\right)}$
-$\leq\frac{2\sqrt{A_{t}}}{L^{3/2}\operatorname{Var}(F_{L}(t))}$
-uniformly for all $t\in[0,T]$ and $L\geq 1$. Proposition 3.1 ensures that
-$\operatorname{Var}(F_{L}(t))\sim\frac{1}{L}\int_{0}^{t}e^{-2\alpha(t-s)}f_{\sigma}(s)\ \mathrm{d}s$
-as $L\to\infty$. This together with (11) implies (4). ∎
+$\leq\frac{2\sqrt{A_{t}}}{L^{3/2}\operatorname{Var}(F_{L}(t))}.$
+From Proposition 3.1, the asymptotic behavior of the variance is given by
+$\operatorname{Var}(F_{L}(t))\gtrsim\frac{1}{L}\frac{\sigma(1)^{2}}{2}\int_{0}^{t\wedge\delta}e^{-2\alpha(t-s)}\ \mathrm{d}s,$
+as $L\to\infty$. Consequently, using this result along with $\sigma(1)\neq 0$ and (19), we can deduce (4). ∎
 5 Proof of Theorem 2
@@
 $\int_{0}^{t_{1}}\int_{0}^{L}(\mathcal{I}_{0}(t_{2}-t_{1}+s,y)-\mathcal{I}_{0}(s,y))^{2}\ \mathrm{d}y\mathrm{d}s$
-$\leq 16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\int_{0}^{t_{1}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\ \mathrm{d}s\left(|\alpha|^{2}e^{2|\alpha|T}|t_{2}-t_{1}|^{2}+\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}\right)$
-$=\frac{16}{6}L|\alpha|^{2}Te^{4|\alpha|T}|t_{2}-t_{1}|^{2}+16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\int_{0}^{t_{1}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\ \mathrm{d}s\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}.$
+[…]
+[…]
 Using the fact that $1-e^{-x}\leq 1\wedge x$ for all $x\geq 0$,
-$16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\int_{0}^{t_{1}}e^{-\frac{n^{2}\pi^{2}s}{L^{2}}}\ \mathrm{d}s\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}$
-$=16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\frac{L^{2}}{n^{2}\pi^{2}}\left(1-e^{-\frac{n^{2}\pi^{2}t_{1}}{L^{2}}}\right)\left|e^{-\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}}-1\right|^{2}$
-$\leq 16Le^{2|\alpha|T}\sum_{n=1}^{\infty}\frac{1}{n^{2}\pi^{2}}\left|1\wedge\frac{n^{2}\pi^{2}(t_{2}-t_{1})}{2L^{2}}\right|^{2}\frac{L^{2}}{n^{2}\pi^{2}}$
+[…]
+[…]
+[…]
 $\leq\frac{16e^{2|\alpha|T}}{L}\sum_{n=1}^{\infty}\left(\frac{L^{4}}{n^{4}\pi^{4}}\wedge|t_{2}-t_{1}|^{2}\right).$
@@
 $\leq\frac{2z_{k}^{2}M_{\sigma}^{2}(1+c_{T,k})^{2}}{L}\left(e^{2|\alpha|T}+\frac{16}{6}|\alpha|^{2}T^{2}e^{4|\alpha|T}+\frac{176}{3\pi}T^{1/2}e^{2|\alpha|T}\right)|t_{2}-t_{1}|.$
-Therefore, in the case of any boundary conditions, we have (12). This completes the proof. ∎
+Therefore, in the case of Neumann, Dirichlet, or periodic boundary conditions, we have (20). This completes the proof. ∎
 Proof of Theorem 2.
-The tightness of $\{\sqrt{L}F_{L}\}_{L\geq 1}$ in $C([0,T])$ is a consequence of Proposition 5.1 (see [K98]).
-Then, it remains to prove that the finite-dimensional distributions of the process $(\sqrt{L}F_{L}(t))_{t\in[0,T]}$ converge to those of $(\int_{0}^{t}e^{-\alpha(t-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s})_{t\in[0,T]}$ as $L\to\infty$.
-Let us fix $T>0$ and $m\geq 1$ points $t_{1},\dots,t_{m}\in(0,T]$. Proposition 3.1 ensures that, for every $i,j=1,\dots,m$,
+The tightness of the family of processes $\{\sqrt{L}F_{L}(\cdot)\}_{L\geq 1}$ in $C([0,T])$ is guaranteed by Proposition 5.1 (see, e.g., [K98]).
+What remains is to show that the finite-dimensional distributions of $(\sqrt{L}F_{L}(t))_{t\in[0,T]}$ converge to those of $(\int_{0}^{t}e^{-\alpha(t-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s})_{t\in[0,T]}$ as $L\to\infty$
+Let us fix $T>0$ and $m\geq 1$ points $t_{1},\dots,t_{m}\in(0,T]$. According to Proposition 3.2, as $L\to\infty$, the covariance satisfies
 $\operatorname{Cov}\left[F_{L}(t_{i}),F_{L}(t_{j})\right]\sim\frac{1}{L}\int_{0}^{t_{i}\wedge t_{j}}e^{-\alpha(t_{i}+t_{j}-2s)}f_{\sigma}(s)\ \mathrm{d}s$ (23)
-as $L\to\infty$. We define the following quantities:
-$F_{i}:=\frac{F_{L}(t_{i})}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}}$
-and
-$C_{i,j}:=\operatorname{Cov}\left[F_{i},F_{j}\right]$
-for $i,j=1,\dots,m$. We will write $F:=(F_{1},\dots,F_{m})$, and let $G=(G_{1},\dots,G_{m})$ denote a centered Gaussian random vector with covariance matrix $C=(C_{i,j})_{1\leq i,j\leq m}$.
-Recall from (9) the random fields $v_{L,t_{1}},\dots,v_{L,t_{m}}$, and define rescaled random fields $V_{1},\dots,V_{m}$ as follows:
-$V_{i}:=\frac{v_{L,t_{i}}}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}}$
-for $i=1,\dots,m$.
-According to (8), $F_{i}=\delta(V_{i})$ for all $i=1,\dots,m$.
-And, by duality, $E[\langle DF_{i},V_{j}\rangle_{\mathcal{H}}]=E[F_{i}\delta(V_{j})]=C_{i,j}$ for all $i,j=1,\dots,m$.
-Then, Proposition 2.2 implies that
-$\left|E[h(F)]-E[h(G)]\right|\leq\frac{1}{2}\|h^{\prime\prime}\|_{\infty}\sqrt{\sum_{i,j=1}^{m}\operatorname{Var}\left(\langle DF_{i},V_{j}\rangle_{\mathcal{H}}\right)}$
-for all $h\in C^{2}(\mathbb{R}^{m})$. From Proposition 4.1,
+for all $i,j=1,\dots,m$. To proceed, let us define the vector $F:=(F_{1},\dots,F_{m})$ with components
+$F_{i}:=\frac{F_{L}(t_{i})}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}}.$
+We then define $G=(G_{1},\dots,G_{m})$ to be a centered Gaussian random vector whose covariance matrix $C=(C_{i,j})$ is given by
+$C_{i,j}:=\operatorname{Cov}\left[F_{i},F_{j}\right].$
+Let us introduce the rescaled random fields $V_{1},\dots,V_{m}$ by setting
+$V_{i}:=\frac{v_{L,t_{i}}}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}},\quad i=1,\dots,m,$
+where the fields $v_{L,t_{i}}$ are given in (17).
+From (16), we have the relation $F_{i}=\delta(V_{i})$.
+By duality, $E[\langle DF_{i},V_{j}\rangle_{\mathcal{H}}]=E[F_{i}\delta(V_{j})]=C_{i,j}$ for all $i,j=1,\dots,m$.
+Applying Proposition 2.2 then yields the desired bound for any $h\in C^{2}(\mathbb{R}^{m})$:
+$\left|E[h(F)]-E[h(G)]\right|\leq\frac{1}{2}\|h^{\prime\prime}\|_{\infty}\sqrt{\sum_{i,j=1}^{m}\operatorname{Var}\left(\langle DF_{i},V_{j}\rangle_{\mathcal{H}}\right)}.$
+From Proposition 4.1,
 $\operatorname{Var}\left(\langle DF_{i},V_{j}\rangle_{\mathcal{H}}\right)$ $=\frac{\operatorname{Var}\left(\langle DF_{L}(t_{i}),v_{L,t_{j}}\rangle_{\mathcal{H}}\right)}{\operatorname{Var}(F_{L}(t_{i}))\operatorname{Var}(F_{L}(t_{j}))}$
@@
 for all $h\in C^{2}(\mathbb{R}^{m})$.
-On the other hand, owing to (15), as $L\to\infty$,
-$C_{i,j}=\frac{\operatorname{Cov}(F_{L}(t_{i}),F_{L}(t_{j}))}{\sqrt{\operatorname{Var}(F_{L}(t_{i}))}\sqrt{\operatorname{Var}(F_{L}(t_{j}))}}\to\frac{\int_{0}^{t_{i}\wedge t_{j}}e^{-\alpha(t_{i}+t_{j}-2s)}f_{\sigma}(s)\ \mathrm{d}s}{\sqrt{\int_{0}^{t_{i}}e^{-2\alpha(t_{i}-s)}f_{\sigma}(s)\ \mathrm{d}s}\sqrt{\int_{0}^{t_{j}}e^{-2\alpha(t_{j}-s)}f_{\sigma}(s)\ \mathrm{d}s}},$
-which yields that the Gaussian random vector $G$ converges weakly to
-$\left(\frac{\int_{0}^{t_{1}}e^{-\alpha(t_{1}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{1}}e^{-2\alpha(t_{1}-s)}f_{\sigma}(s)\ \mathrm{d}s}},\dots,\frac{\int_{0}^{t_{m}}e^{-\alpha(t_{m}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{m}}e^{-2\alpha(t_{m}-s)}f_{\sigma}(s)\ \mathrm{d}s}}\right)$ (16)
-as $L\to\infty$.
-Therefore, it follows that $F$ converges weakly to the random vector in (16) as $L\to\infty$. Then, from (15), it follows that
+We also examine the limit of the covariances $C_{i,j}$. Using the result from (23), we find that as $L\to\infty$,
+$C_{i,j}\to\frac{\int_{0}^{t_{i}\wedge t_{j}}e^{-\alpha(t_{i}+t_{j}-2s)}f_{\sigma}(s)\ \mathrm{d}s}{\sqrt{\int_{0}^{t_{i}}e^{-2\alpha(t_{i}-s)}f_{\sigma}(s)\ \mathrm{d}s}\sqrt{\int_{0}^{t_{j}}e^{-2\alpha(t_{j}-s)}f_{\sigma}(s)\ \mathrm{d}s}}.$
+It follows that the law of the Gaussian vector $G$, which is determined by its covariance structure, converges weakly to that of
+$\left(\frac{\int_{0}^{t_{1}}e^{-\alpha(t_{1}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{1}}e^{-2\alpha(t_{1}-s)}f_{\sigma}(s)\ \mathrm{d}s}},\dots,\frac{\int_{0}^{t_{m}}e^{-\alpha(t_{m}-s)}\sqrt{f_{\sigma}(s)}\ \mathrm{d}W_{s}}{\sqrt{\int_{0}^{t_{m}}e^{-2\alpha(t_{m}-s)}f_{\sigma}(s)\ \mathrm{d}s}}\right).$ (24)
+Therefore, $F$ converges weakly to the random vector in (24) as $L\to\infty$. Recalling the definition of $F_{i}$ and using the asymptotic variance from (23), the random vector
 $\sqrt{L}\left(\frac{F_{L}(t_{1})}{\sqrt{\int_{0}^{t_{1}}e^{-2\alpha(t_{1}-s)}f_{\sigma}(s)\ \mathrm{d}s}},\dots,\frac{F_{L}(t_{m})}{\sqrt{\int_{0}^{t_{m}}e^{-2\alpha(t_{m}-s)}f_{\sigma}(s)\ \mathrm{d}s}}\right)$
@@
 and
-$\mathcal{I}_{k}(t,x):=$ $\sigma_{1}^{k-1}\int_{0}^{t}\int_{0}^{L}\dots\int_{0}^{r_{k-1}}\int_{0}^{L}G_{t-r_{1}}(x,z_{1})\dots G_{r_{k-1}-r_{k}}(z_{k-1},z_{k})$
-×$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})\ W(\mathrm{d}r_{k},\mathrm{d}z_{k})\dots W(\mathrm{d}r_{1},\mathrm{d}z_{1})$
+$\mathcal{I}_{k}(t,x)$= $\sigma_{1}^{k-1}\int_{0}^{t}\int_{0}^{L}\dots\int_{0}^{r_{k-1}}\int_{0}^{L}G_{t-r_{1}}(x,z_{1})\dots G_{r_{k-1}-r_{k}}(z_{k-1},z_{k})$
+$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})\ W(\mathrm{d}r_{k},\mathrm{d}z_{k})\dots W(\mathrm{d}r_{1},\mathrm{d}z_{1})$
 for all $k\geq 1$. We define the Picard iteration $\{u_{n}(t,x)\}_{n=0}^{\infty}$ for $u(t,x)$. Define iteratively, for every $n\geq 0$,
@@
 $\|\mathcal{I}_{k}(t,x)\|_{2}^{2}=$ $\sigma_{1}^{2(k-1)}\int_{0}^{t}\int_{0}^{L}\dots\int_{0}^{r_{k-1}}\int_{0}^{L}G_{t-r_{1}}^{2}(x,z_{1})\dots G_{r_{k-1}-r_{k}}^{2}(z_{k-1},z_{k})$
-×$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
+$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
 for all $k\geq 1$.
@@
 $\leq\sigma_{1}^{2(k-1)}K_{T}^{2k}(|\sigma_{1}|e^{|\alpha|T}+|\sigma_{0}|)^{2}$
-×$\quad\times\int_{0}^{t}\int_{\mathbb{R}}\dots\int_{0}^{r_{k-1}}\int_{\mathbb{R}}p_{\beta(t-r_{1})}^{2}(x-z_{1})\dots p_{\beta(r_{k-1}-r_{k})}^{2}(z_{k-1}-z_{k})\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
+$\quad\times\int_{0}^{t}\int_{\mathbb{R}}\dots\int_{0}^{r_{k-1}}\int_{\mathbb{R}}p_{\beta(t-r_{1})}^{2}(x-z_{1})\dots p_{\beta(r_{k-1}-r_{k})}^{2}(z_{k-1}-z_{k})\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
 $=\sigma_{1}^{2(k-1)}K_{T}^{2k}(|\sigma_{1}|e^{|\alpha|T}+|\sigma_{0}|)^{2}\left(\frac{t}{4\pi\beta}\right)^{k/2}$
-×$\quad\times\int_{0<r_{k}<\dots<r_{1}<1}\sqrt{\frac{1}{(1-r_{1})\times\dots\times(r_{k-1}-r_{k})}}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
+$\quad\times\int_{0<r_{k}<\dots<r_{1}<1}\sqrt{\frac{1}{(1-r_{1})\times\dots\times(r_{k-1}-r_{k})}}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
 $=\sigma_{1}^{2(k-1)}K_{T}^{2k}(|\sigma_{1}|e^{|\alpha|T}+|\sigma_{0}|)^{2}\left(\frac{t}{4\pi\beta}\right)^{k/2}\frac{\Gamma(1/2)^{k}}{\Gamma((k+2)/2)}$
@@
 $=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}e^{-2\alpha(t-r_{1})}p_{2\beta(t-r_{1})}(0)\dots e^{-2\alpha(r_{k-1}-r_{k})}p_{2\beta(r_{k-1}-r_{k})}(0)$
-×$\quad\times\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
+$\quad\times\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
 $\quad+o(L).$
@@
 $=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}e^{-2\alpha(t-r_{1})}p_{2\beta(t-r_{1})}(0)\dots e^{-2\alpha(r_{k-1}-r_{k})}p_{2\beta(r_{k-1}-r_{k})}(0)$
-×$\quad\times\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
+$\quad\times\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
 $=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}e^{-2\alpha(t-r_{1})}p_{2\beta(t-r_{1})}(0)\dots e^{-2\alpha(r_{k-1}-r_{k})}p_{2\beta(r_{k-1}-r_{k})}(0)$
-×$\quad\times(\sigma_{1}^{2}e^{-2\alpha r_{k}}+2\sigma_{1}\sigma_{0}e^{-\alpha r_{k}}+\sigma_{0}^{2})\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
+$\quad\times(\sigma_{1}^{2}e^{-2\alpha r_{k}}+2\sigma_{1}\sigma_{0}e^{-\alpha r_{k}}+\sigma_{0}^{2})\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
 $=\sigma_{1}^{2(k-1)}e^{-2\alpha t}\int_{0<r_{k}<\dots<r_{1}<t}p_{2\beta(t-r_{1})}(0)\dots p_{2\beta(r_{k-1}-r_{k})}(0)$
-×$\quad\times(\sigma_{1}+e^{\alpha r_{k}}\sigma_{0})^{2}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
+$\quad\times(\sigma_{1}+e^{\alpha r_{k}}\sigma_{0})^{2}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
 $=:f_{k}(t).$
@@
 $=\sigma_{1}^{2(k-1)}\int_{0}^{t}\int_{0}^{L}\dots\int_{0}^{r_{k-1}}\int_{0}^{L}G_{t-r_{1}}^{2}(x,z_{1})\dots G_{r_{k-1}-r_{k}}^{2}(z_{k-1},z_{k})$
-×$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
+$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
 $=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}G_{2(t-r_{1})}(0,0)\dots G_{2(r_{k-1}-r_{k})}(0,0)(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
@@
 $=\sigma_{1}^{2(k-1)}\int_{0<r_{k}<\dots<r_{1}<t}e^{-2\alpha(t-r_{1})}p_{2\beta(t-r_{1})}(0)\dots e^{-2\alpha(r_{k-1}-r_{k})}p_{2\beta(r_{k-1}-r_{k})}(0)$
-×$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
+$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
 $=\sigma_{1}^{2(k-1)}e^{-2\alpha t}\int_{0<r_{k}<\dots<r_{1}<t}p_{2\beta(t-r_{1})}(0)\dots p_{2\beta(r_{k-1}-r_{k})}(0)$
-×$\quad\times(\sigma_{1}+\sigma_{0}e^{\alpha r_{k}})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
+$\quad\times(\sigma_{1}+\sigma_{0}e^{\alpha r_{k}})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
 $=f_{k}(t),$
@@
 where in the second equality, we apply dominated convergence theorem in order to exchange the limit and the sum.
-Therefore, in the case of any boundary conditions, we obtain
+Therefore, in the case of Neumann, Dirichlet, or periodic boundary conditions, we obtain
 $\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}E[u(t,x)^{2}]\ \mathrm{d}x=\sum_{k=0}^{\infty}f_{k}(t).$
@@
 $\sigma_{1}^{2}f_{k}(t)$ $\leq\sigma_{1}^{2k}e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}$
-×$\quad\times\int_{0<r_{k}<\dots<r_{1}<t}p_{2\beta(t-r_{1})}(0)\dots p_{2\beta(r_{k-1}-r_{k})}(0)\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
+$\quad\times\int_{0<r_{k}<\dots<r_{1}<t}p_{2\beta(t-r_{1})}(0)\dots p_{2\beta(r_{k-1}-r_{k})}(0)\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
 $=e^{-2\alpha t}(|\sigma_{1}|+e^{|\alpha|t}|\sigma_{0}|)^{2}\frac{(\sigma_{1}^{4}t/(4\beta))^{k/2}}{\Gamma((k+2)/2)}.$
@@
 Hence, the properties of $G$ follow directly from those of $g$. For all $t>0$ and $x,y\in[0,L]$, in the case of Neumann boundary conditions,
-$G_{t}(x,y)=e^{-\alpha t}\sum_{n\in\mathbb{Z}}\left(p_{\beta t}(x-y+2nL)+p_{\beta t}(x+y+2nL)\right),$
+$G_{t}(x,y)=e^{-\alpha t}\sum_{n\in\mathbb{Z}}\left(p_{\beta t}(x-y+2nL)+p_{\beta t}(x+y+2nL)\right),$ (35)
 or equivalently,
@@
 and in the case of periodic boundary conditions,
-$G_{t}(x,y)=e^{-\alpha t}\sum_{nin\mathbb{Z}}p_{\beta t}(x-y+nL).$
+$G_{t}(x,y)=e^{-\alpha t}\sum_{n\in\mathbb{Z}}p_{\beta t}(x-y+nL).$ (37)
 Moreover, by using the properties of $g$ [P22]*Lemma A.1., we get some useful properties of $G$.